### v1

# Typical Quantum States of the Universe are Observationally Indistinguishable

Abstract: This paper is about the epistemology of quantum theory. We establish a new result about a limitation to knowledge of its central object—the quantum state of the universe. We show that, if the universal quantum state can be assumed to be a typical unit vector from a high-dimensional subspace of Hilbert space (such as the subspace defined by a low-entropy macro-state as prescribed by the Past Hypothesis), then no observation can determine (or even just narrow down significantly) which vector it is. Typical state vectors, in other words, are observationally indistinguishable from each other. Our argument is based on a typicality theorem from quantum statistical mechanics. We also discuss how theoretical considerations that go beyond the empirical evidence might bear on this fact and on our knowledge of the universal quantum state.
Key words: limitation to knowledge; empirical equivalence; past hypothesis; quantum statistical mechanics.
1 Introduction
This paper is about the epistemology of quantum theory. We establish a new result about a limitation to knowledge of the quantum state of the universe. We show that typical universal quantum states are observationally indistinguishable, in the sense that no observation can distinguish typical universal quantum states in a high-dimensional subspace of the Hilbert space, such as the subspace characterized by a low-entropy macro-state as prescribed by the Past Hypothesis (see §2.2 and (Albert 2000, Goldstein et al. 2020)). We call this fact observation typicality.
Observation typicality follows from a result in quantum statistical mechanics called distribution typicality (Reimann 2007, Teufel et al. 2023): for any observable, most quantum states in a high-dimensional Hilbert subspace lead to very nearly the same probability distribution for that observable. We show that this implies observation typicality.
Observation typicality is a surprisingly strong result. There are other known examples of epistemic limitations in physics. Dürr et al. (2004) already proved that it is impossible to measure a wave function precisely. Observation typicality is a further step in this direction, and it delivers a stronger conclusion—even an imprecise measurement is impossible. No observation, as long as it does not disconfirm the physical theory, will give us any substantive information about the quantum state. In this sense, limitation to knowledge is more pervasive and persistent in a quantum universe than has been recognized.
A statement about most quantum states in the full Hilbert space $\mathscr{H}$ of the universe has limited applicability because it may fail for the quantum states of interest. For example, most quantum states in $\mathscr{H}$ are in thermal equilibrium, but the actual quantum state of our universe is not. To use typicality theorems in our situation, we apply them to a subspace $\mathscr{H}_{0}$ of $\mathscr{H}$ such as the subspace comprising the quantum states $\Psi$ compatible with a certain macro-state of the universe that may be far from thermal equilibrium. We assume that $\mathscr{H}_{0}$ has finite but huge dimension $d_{0}$ (while $\mathscr{H}$ may have finite or infinite dimension). Observation typicality then holds for most quantum states $\Psi$ in $\mathscr{H}_{0}$. We can use the result while accepting certain inductive hypotheses about the physical laws of the universe, including those pertaining to initial conditions, according to which our universe is a typical member of the possible universes compatible with such laws.
We assume here that there is a wave function $\Psi_{t}$ of the universe and that at the initial time $t_{0}$ of the universe (say, at the big bang), $\Psi_{t_{0}}$ had to lie in a particular subspace $\mathscr{H}_{0}$ of the Hilbert space $\mathscr{H}$ of the universe (say, corresponding to a low-entropy macro-state, as prescribed by the Past Hypothesis). Observation typicality asserts that our empirical data at any time $t$ will reveal very little about $\Psi_{t_{0}}$ (beyond its membership in $\mathscr{H}_{0}$, which we take as known) because most $\Psi_{t_{0}}$ from $\mathscr{H}_{0}$ lead to very nearly the same probability distribution for these empirical data. In other words, typical initial quantum states of the universe are observationally indistinguishable from each other. If the time evolution is unitary, then the same arguments yield that also $\Psi_{t}$ cannot be distinguished empirically from typical alternative vectors in the appropriate subspace, i.e., in $U_{t}\mathscr{H}_{0}$ with $U_{t}=\exp(-iH(t-t_{0}))$ and $H$ the Hamiltonian of the universe.
The analogous statement fails in classical mechanics. Observation typicality holds in quantum mechanics because, roughly speaking, quantum states can be in superpositions, and the superposition weights of any particular observation given by typical quantum states are nearly the same. The latter is related to the mathematical fact, known as the concentration of measure phenomenon (the situation where random quantities become nearly deterministic), which applies in any high-dimensional Hilbert space.
We present two main results. First (§2 and §3), that a typical unit vector $\Psi_{t_{0}}\in\mathscr{H}_{0}$ cannot be reliably distinguished from the density matrix $\rho_{0}$ associated with a uniform distribution in $\mathscr{H}_{0}$, i.e., $\rho_{0}=P_{0}/d_{0}$ with $P_{0}$ the projection operator to $\mathscr{H}_{0}$ and $d_{0}=\dim\mathscr{H}_{0}$; as a consequence, typical unit vectors cannot be reliably distinguished from each other. The second result (§3.4) is even stronger and asserts that even unreliably, they cannot be distinguished; that is, our empirical data do not even yield (any significant amount of) partial or probabilistic information about $\Psi_{t_{0}}$. This result can be expressed in terms of Bayesian credences as follows: if we start from the uniform probability distribution $u_{0}$ over the unit sphere $\mathbb{S}(\mathscr{H}_{0})$ (containing all normalized wave functions) in $\mathscr{H}_{0}$ as the prior distribution, and if we update our credences in a Bayesian way on the basis of our empirical observation (that is not too improbable), then the updated distribution will still be very nearly uniform over $\mathbb{S}(\mathscr{H}_{0})$.
Some empirical observations might, of course, change our mind about which subspace $\mathscr{H}_{0}$ is the correct one. Since we regard here the specification of $\mathscr{H}_{0}$ as part of the given theory, such an observation would amount to a disconfirmation of the theory. This leads to the interesting question of which kind of observations should be regarded as disconfirming the theory. We will look more into this question elsewhere (Chen and Tumulka 2024), take here for granted that the theory is correct and not getting disconfirmed, and remark only briefly that since the specification of $\mathscr{H}_{0}$ has the status of a law of nature, we may expect it to be simple and natural in order to be convincing; as a consequence, we would not be inclined to adjust the choice of $\mathscr{H}_{0}$ in complicated ways just to adapt it to some random-looking features of our world (such as, e.g., the shape of Ireland).
Our goal is to further the discussion about the epistemology of physical theories and highlight the limitations of (direct) empirical knowledge in quantum universes. Insofar as we have detailed knowledge about the universal quantum state, it is even more theoretical than has been recognized. Our result should be of interest to anyone interested in the epistemological implications of physics. We highlight several distinctive features of the present approach:
  • Observation typicality applies to all interpretations of quantum mechanics. It is also compatible with generous inductive assumptions about what we can infer based on available evidence. For example, it is compatible with our knowledge of the dynamical laws of quantum mechanics, the Past Hypothesis, and the aggregation of individual observations into a collection.
  • Our argument relies on the concentration of measure phenomenon instead of special constructive techniques, such as the cut-and-paste method used in Manchak (2009)’s argument for observational indistinguishability in general relativity. The standard “physical unreasonableness” objections against arguments based on special constructions do not apply to our argument, which establishes observational indistinguishability for generic pairs of models.
  • Our claims are based on quantitative bounds (recently proven mathematically) on how large the deviations of certain probabilities can at most become for which fraction of all wave functions in a high-dimensional Hilbert subspace.
In §2, we formulate distribution typicality and explain how it applies to the special case of the Past Hypothesis. In §3, we clarify our criterion for observational indistinguishability, derive observation typicality from distribution typicality, apply the fact to the universal quantum state, and provide a Bayesian analysis. In §4, we compare observation typicality to known results about epistemic limitations in physics. In §5, we discuss potential philosophical implications. For more detail, see the longer version, Chen and Tumulka (2024).
2 Distribution Typicality
In the 21st century, many exciting new results have been proven in quantum statistical mechanics and quantum thermodynamics, leading to improved understanding in the thermodynamic behavior of a closed quantum system in a pure state (Gemmer et al. 2009, Tasaki 2016, Mori et al. 2018). Such results often assume an “individualist” attitude, according to which an individual quantum system in a pure state, as opposed to an ensemble, is studied with respect to its thermodynamic properties. These results often concern typical properties of those states. [^3] There is a growing philosophical literature about typicality and its role in physical theories. See, for example, Goldstein (2012), Reichert (2018), Wilhelm (2022), Lazarovici (2023), Hubert (2024), (Frigg and Werndl 2024, sec. 4.5), and the references therein.
We shall focus on distribution typicality in this paper. In §3, we show that distribution typicality implies observation typicality. In this section, we formulate distribution typicality and specialize it to the Past Hypothesis.
2.1 The Key Theorem
Before stating the key theorem, we highlight its implications for orthodox quantum mechanics (OQM), Bohmian mechanics (BM), and the Ghirardi-Rimini-Weber spontaneous collapse theory (GRW) (Tumulka 2022b, Dürr and Teufel 2009, Ghirardi and Bassi 2020). Distribution typicality entails that, in OQM, if an observer at time $t$ conducts an experiment, the probability distribution of the outcome is nearly independent of $\Psi_{t_{0}}$; in BM, at any time $t$, the probability distribution of the Bohmian configurations $Q_{t}$ of the universe is nearly independent of $\Psi_{t_{0}}$; in GRW with a flash ontology, for any time $t>t_{0}$, the probability distribution of the pattern of flashes up to $t$ is nearly independent of $\Psi_{t_{0}}$. We will elucidate in §3 why this near-independence entails the impossibility for observers according to OQM, BM, or GRW to distinguish among different $\Psi_{t_{0}}$.
Terminology. Recall that $\mathbb{S}(\mathscr{H}_{0})$ is the unit sphere in the subspace $\mathscr{H}_{0}$, $d_{0}=\dim\mathscr{H}_{0}$, and $u_{0}$ the uniform probability distribution over $\mathbb{S}(\mathscr{H}_{0})$; that is, for any subset $S$ of $\mathbb{S}(\mathscr{H}_{0})$, $u_{0}(S)$ is the surface area of $S$, normalized by dividing through the surface area of $\mathbb{S}(\mathscr{H}_{0})$. We say that a statement $s(\Psi)$ is true for $(1-\varepsilon)$-most $\Psi\in\mathbb{S}(\mathscr{H}_{0})$ if and only if the set $S$ of $\Psi\in\mathbb{S}(\mathscr{H}_{0})$ satisfying $s(\Psi)$ has $u_{0}(S)\geq 1-\varepsilon$ (equivalently, if $S$ has at least the fraction $1-\varepsilon$ of the surface area, or if a purely random point on the sphere has probability $\geq 1-\varepsilon$ to lie in $S$). Averages on $\mathbb{S}(\mathscr{H}_{0})$ will also be taken with respect to $u_{0}$. [^4] The results in this paper can be extended to the Gaussian adjusted projected (GAP) measures introduced by Goldstein et al. (2006b), as is done by Teufel et al. (2024). But here, we consider only the uniform measure $u_{0}$. We also need the concept of POVM.
POVM. POVM stands for positive-operator-valued measure, which is a generalization of the concept of an observable as given by a self-adjoint operator. Technically, a POVM on a Hilbert space $\mathscr{H}$ is a family of positive operators $E_{z}$ on $\mathscr{H}$ that add up to the identity operator. POVMs have an important representational role in quantum mechanics, given by the Main Theorem about POVMs: for every quantum physical experiment $\mathscr{E}$ on a quantum system $\mathscr{S}$ whose possible outcomes lie in a space $\mathscr{Z}$, there exists a POVM E on $\mathscr{Z}$ such that, whenever $\mathscr{S}$ has wave function $\Psi$ at the beginning of $\mathscr{E}$, the random outcome $Z$ has probability distribution given by
$\mathbb{P}(Z=z)=\langle\Psi|\,E_{z}\,|\Psi\rangle.$ (1)
See Dürr et al. (2004) and Tumulka (2022b) for proofs.
The Main Theorem about POVMs provides a generalization of the Born rule to arbitrary experiments (instead of ideal quantum measurements). The operator $E_{z}$ is associated with the possible outcome $z$; in the case of an ideal quantum measurement, $E_{z}$ would be the projection operator to the eigenspace of the eigenvalue $z$.
We can now state the key theorem as follows:
Theorem 1 (Distribution Typicality).
Let $\mathscr{H}_{0}$ be an arbitrary subspace with finite dimension $d_{0}$ in $\mathscr{H}$, let the density matrix $\rho_{0}=P_{0}/d_{0}$ be the normalized projection to $\mathscr{H}_{0}$, and let the operator $E_{z}$ be an element of an arbitrary POVM on $\mathscr{H}$. Then for every $\varepsilon>0$, for $(1-\varepsilon)$-most $\Psi\in\mathbb{S}(\mathscr{H}_{0})$,
$\Bigl{|}\langle\Psi|\,E_{z}\,|\Psi\rangle-\operatorname{tr}(\rho_{0}\,E_{z})\Bigr{|}\leq\frac{1}{\sqrt{\varepsilon d_{0}}}\,.$ (2)
Theorem 1 is a particular case of Theorem 3 proven by Teufel et al. (2023). It can also be derived from a theorem proven by Reimann (2007).
It can be shown that $\operatorname{tr}(\rho_{0}E_{z})$ is exactly the average of $\langle\Psi|E_{z}|\Psi\rangle$ over $\mathbb{S}(\mathscr{H}_{0})$ (using the uniform distribution $u_{0}$); thus (2) expresses that $\langle\Psi|E_{z}|\Psi\rangle$ is close to its average value, provided that the right-hand side is small, which will be the case whenever
$d_{0}\gg\frac{1}{\varepsilon}\,.$ (3)
Inequality (2) is an example of the “concentration of measure phenomenon,” which can be expressed as saying that some random quantity is nearly deterministic (i.e., very probably close to a deterministic value). Here, we have that for any observation represented by an element in a POVM, typical quantum states in a high-dimensional Hilbert space have very nearly the same Born probability distribution for that observation.
Let us clarify and elaborate on this result.
Example. A macroscopic system with $N=10^{20}$ particles is in macro-state $\mu$. The dimension of the Hilbert space $\mathscr{H}_{0}$ of the system scales exponentially with $N$. Let us suppose $d_{0}=10^{10^{20}}$. Consider a very small error $\varepsilon=10^{-200}$. Inequality (2) tells us that, for any quantum experiment with observational outcome $z$, for $(1-\varepsilon)$-most pure states in $\mathscr{H}_{0}$, the Born probability of outcome $z$ given by the pure state differs from the ensemble average by at most $10^{100-0.5\times 10^{20}}\approx 3^{-10^{20}}$, which is close to zero.
Minimal Assumptions. Theorem 1 holds in great generality. First, it applies to every POVM. By the Main Theorem about POVMs, the theorem covers every observational outcome arising from an arbitrary quantum physical experiment. Second, it does not make assumptions about the interaction between the subsystem and the environment. Third, it does not invoke chaos, ergodicity, or mixing. The typicality in distribution typicality comes from the typicality of quantum states in $\mathbb{S}(\mathscr{H}_{0})$, and the large number comes from $d_{0}$, the large dimension of $\mathscr{H}_{0}$. Finally, distribution typicality applies to thermal equilibrium and non-equilibrium. For concreteness, in §2.2 we provide a physical interpretation for a universe in thermal non-equilibrium, taking $\mathscr{H}_{0}$ to be the low-entropy initial macro-state of the universe prescribed by the Past Hypothesis.
Difference from the Classical Case. We do not have distribution typicality in classical mechanics. The prediction of any observation from a particular micro-state in the classical phase space has a trivial probability of 00 or $1$, while the expectation values for many observables with respect to a uniform probability distribution (the ensemble averages) are often non-trivial. It is false that for every experimental outcome, typical micro-states will assign very nearly the same probability distribution. For example, we can design a coin-flip experiment where half the microstates assign probability 00 and the other half assign probability $1$ to the “Tails” outcome. The relevant way in which quantum mechanics differs from classical mechanics is that if a wave function $\Psi=\Psi_{t_{0}}$ gets chosen randomly, then the outcome $Z$ of an observation is “doubly random”: Given $\Psi$, it is random with the Born distribution $\langle\Psi|E_{z}|\Psi\rangle$, and in addition $\Psi$ is itself random. The crucial point is that in our situation, different $\Psi$s have very similar superposition weights, with the consequence that they have nearly the same $\langle\Psi|E_{z}|\Psi\rangle$.
2.2 Main Motivation: The Past Hypothesis
A typicality statement about most quantum states in some space has limited applicability if the relevant quantum states are atypical in that space. For example, most quantum states in the energy shell are in thermal equilibrium, but many quantum states we observe are not. That is not a problem in our case, as Theorem 1 is quite general. To apply Theorem 1 to the relevant class of quantum states, we can specialize the arbitrary subspace $\mathscr{H}_{0}$ to any macro-state that may be far from thermal equilibrium. We are particularly motivated by the thought that the physical laws may require the initial wave function of the universe to lie in a particular subspace $\mathscr{H}_{0}$.
In fact, it has been suggested for the explanation of the thermodynamic arrow of time (Feynman 1963, p. 115), (Penrose 1979, Lebowitz 1993, Albert 2000, Goldstein et al. 2020) that the initial state of the universe must be restricted to a set of low-entropy states; in a quantum theory, such a set would be given by a suitable subspace $\mathscr{H}_{0}$ of the Hilbert space $\mathscr{H}$ of the universe, and the presumed additional law might then be formulated as follows:
Past Hypothesis (PH)
$\Psi_{t_{0}}$ is a typical element of $\mathscr{H}_{0}$.
Here, $\mathscr{H}_{0}$ is assumed to contain all wave functions compatible with a certain macro-state. For example, Penrose’s Weyl curvature hypothesis might amount to taking as $\mathscr{H}_{0}$ something like the joint eigenspace with all eigenvalues 0 of all Weyl curvature operators at $t_{0}$. To explain the arrow of time, the initial macro-state should have very low entropy. Taking as the definition of entropy the quantum Boltzmann entropy (Lebowitz 1993, Goldstein et al. 2010, 2020)
$S_{qB}(\Psi_{t_{0}})=k_{B}\text{log}(d_{0})\,,$ (4)
the condition of low entropy corresponds the condition that the dimension $d_{0}$ of $\mathscr{H}_{0}$ is much smaller than the dimension of the full Hilbert space $\mathscr{H}$, in fact much smaller than the dimension of the subspace $\mathscr{H}_{eq}$ corresponding to thermal equilibrium in the same energy shell, $d_{0}\ll d_{eq}$. Then the PH requires $\Psi_{t_{0}}$ to have low entropy and to be far from thermal equilibrium.
Even though $\mathscr{H}_{0}$ has much lower dimension than $\mathscr{H}_{eq}$, $\mathscr{H}_{0}$ is still, like all subspaces representing macro-states, a high-dimensional subspace. Realistically, we can expect $d_{0}>>10^{10^{20}}$. With respect to the normalized uniform measure $u_{0}$, we expect that typical wave functions in $\mathbb{S}(\mathscr{H}_{0})$ will evolve in a way such that $S_{qB}$ increases in the medium and the long run, satisfying the Developmental Conjecture formulated by Goldstein et al. (2020), a version of the Second Law of Thermodynamics for a quantum universe.
Hence, if we choose (say) $\varepsilon=10^{-200}$, then for any $E_{z}$, typical individual wave functions in $\mathscr{H}_{0}$ will all lead to nearly the same probability distribution as $\rho_{0}$. In the philosophical literature, $\rho_{0}$ has been called the Wentaculus density matrix, corresponding to the initial state of the Wentaculus theory (Chen 2021, 2024a).
3 Observation Typicality
In this section, we show that distribution typicality implies observation typicality.
3.1 General Distinctions
There are different kinds of observational indistinguishability (OI) discussed in the philosophical literature. We draw some general distinctions regarding (1) OI and empirical equivalence, and (2) in-principle and in-practice OI.
Empirical equivalence is a relation holding among certain physical theories. Two theories may be regarded as empirically equivalent when they make the same predictions about all possible observations. For example, one can explore whether BM is empirically equivalent to OQM, GRW, or Everettian quantum mechanics by checking whether BM makes the same empirical predictions as the latter.
In contrast, OI is primarily a relation holding among models in the same theory. In the philosophical literature, there is a large literature about the OI of certain models of general relativity. Manchak (2009) showed that for almost every general relativistic spacetime, we can construct another one in such a way that no observation will distinguish between the two spacetimes.
In this paper, we focus on OI as it is applied to models of quantum theory. Here, we shall take an interpretation-neutral approach, and discuss how to apply the result in different interpretations in §3.2. Nevertheless, what we show may also be relevant to the issue of empirical equivalence among quantum theories.
The second distinction is between in-principle OI and in-practice OI. People often focus on exact equivalence of predictions (or retrodictions) about outcomes of observations. This is done in the general relativistic case in terms of exact agreement (up to an isometry transformation) of past lightcones of observers, and in quantum foundations in terms of exact equality of theoretical distributions. In many cases, exact equivalence is sufficient but not necessary for OI. We never did and never will probe the exact microscopic detail of the past lightcone. In practice we will quickly exhaust our resources long before reaching the in-principle limits. Nevertheless, to prove sharp mathematical results, it is sometimes easier to focus on in-principle limits. Of course, in-principle OI implies in-practice OI.
Observation typicality concerns the stronger sense of OI—in principle and not just in practice. We show that in-principle OI follows from approximate equality of probabilistic predictions. The approximate equality of predictions is to be distinguished from the approximation of frequency to chance. For example in Bohmian mechanics, we justify the quantum equilibrium hypothesis using a law-of-large-number argument (Dürr et al. 1992), with frequency converging to chance: typically (with respect to the quantum equilibrium distribution), the empirical distribution converges to the theoretical distribution. The result establishes exact equality of the chance functions (theoretical distributions). While exact equality of probabilities is sufficient for in-principle OI, our argument below shows that it is not necessary.
According to a standard criterion (Nielsen and Chuang 2010, p.86), two quantum states are perfectly distinguishable by some experiment if and only if they are orthogonal. If we do not insist on perfect distinguishability, we may say that two quantum states are reliably distinguishable by some experiment if and only if they are (at least) approximately orthogonal.
We can specialize the criterion for OI to a particular observation $E_{z}$. For two quantum states $\Psi_{1}$ and $\Psi_{2}$, we say that they are perfectly distinguishable by observing the outcome $Z=z$ if and only if $\langle\Psi_{1}|\,E_{z}\,|\Psi_{1}\rangle=0$ and $\langle\Psi_{2}|\,E_{z}\,|\Psi_{2}\rangle=1$ or vice versa; let us assume the former is the case. If we know that the actual quantum state is either $\Psi_{1}$ or $\Psi_{2}$, the outcome $Z=z$ is a perfect indicator of $\Psi_{2}$. The indicator is perfect in the sense that, in 100% of the worlds compatible with the outcome $Z=z$, it correctly indicates the actual quantum state. If we do not insist on perfect distinguishability, we can allow the probabilities to be approximately zero or one, so that they are reliably distinguishable by $E_{z}$ just in case $\langle\Psi_{1}|\,E_{z}\,|\Psi_{1}\rangle\approx 0$ and $\langle\Psi_{2}|\,E_{z}\,|\Psi_{2}\rangle\approx 1$, or $\langle\Psi_{1}|\,E_{z}\,|\Psi_{1}\rangle\approx 1$ and $\langle\Psi_{2}|\,E_{z}\,|\Psi_{2}\rangle\approx 0$, i.e., the absolute difference in their probabilities is close to 1. The outcome $Z=z$ is a reliable indicator of the quantum state in the sense that, in nearly 100% of the worlds compatible with the outcome $Z=z$, it correctly indicates the actual quantum state. This suggests the following sufficient condition:
OI with respect to $E_{z}$
If the Born probability distributions of $E_{z}$ assigned by a set of quantum states are all within (say) $10^{-200}$ of each other, then observing the outcome $Z=z$ does not reliably distinguish among those states.
In what follows, unless specified otherwise, we shall understand OI with this criterion.
3.2 Observational Indistinguishability of Typical Quantum States
In this section, we show that distribution typicality implies observation typicality. We do this separately for OQM, BM, and GRW. The same arguments as for OQM and BM also provide OI for Everettian quantum mechanics.
(1) OI in OQM. Suppose we carry out any experiment $\mathscr{E}$; suppose $\mathscr{E}$ is associated, according to the Main Theorem about POVMs, with the POVM $E$; suppose we obtained the outcome $Z=z$ associated with the operator $E_{z}$. Then typical quantum states in a high-dimensional Hilbert space $\mathscr{H}_{0}$ (with $d_{0}>10^{1000}$) are not distinguished by this outcome from $\rho_{0}$.
Indeed, consider $\mathscr{H}_{0}$ with $d_{0}>10^{1000}$ and $\varepsilon=10^{-200}$. By Theorem 1 for any POVM element $E_{z}$, for ($1-\varepsilon)$-most $\Psi$ in $\mathscr{H}_{0}$, $|\langle\Psi|\,E_{z}\,|\Psi\rangle-\operatorname{tr}(\rho_{0}E_{z})|\leq\frac{1}{\sqrt{\varepsilon d_{0}}}<10^{-400}$. Hence, the Born probability distributions of $E_{z}$ assigned by ($1-\varepsilon)$-most $\Psi$ are all within $\tfrac{1}{2}\varepsilon$ of that of $\rho_{0}$, and thus within $\varepsilon$ of each other. As discussed above, a reliable distinction would require the observation of an event that has probability near 1 in one case and near 0 in the other, which does not happen here. Therefore, observing $E_{z}$ does not reliably distinguish those states from each other or from the normalized projection. So, we arrive at observation typicality in $\mathscr{H}_{0}$.
Again, we may think of $\mathscr{H}_{0}$ as the subspace demanded by the Past Hypothesis. Thus, for any observation $E_{z}$, typical quantum states in $\mathscr{H}_{0}$ are not distinguished by $E_{z}$ from the Wentaculus density matrix. Neither does it distinguish typical quantum states in $\mathscr{H}_{0}$ from each other.
In quantum theory with unitary evolution, this extends to later states too: Since typical $\Psi_{t}$ from the appropriately evolved subspace $U_{t}\mathscr{H}_{0}$ lead to nearly the same probability for outcome $Z=z$, the observation of this outcome does not distinguish among different $\Psi_{t}$.
(2) OI in BM. In BM, a key fact about OI is this:
the probability distribution of the configuration $Q_{t}$ of the universe at time $t$, i.e., the $|\Psi_{t}|^{2}$ distribution in configuration space, is nearly independent of $\Psi_{t_{0}}$ for typical $\Psi_{t_{0}}\in\mathbb{S}(\mathscr{H}_{0})$. (5)
This follows from Theorem 1 by taking the POVM $E$ to be the Heisenberg-evolved configuration observable, i.e., the projection-valued measure jointly diagonalizing all position operators at time $t$, or
$E(B)=U_{t}^{\dagger}1_{B}U_{t},$ (6)
where $B$ is any subset of configuration space and $1_{B}$ the multiplication operator multiplying by the characteristic function of $B$. Note that the probability distribution associated with this POVM is exactly the $|\Psi_{t}|^{2}$ distribution:
[…]
for any subset $B$ of configuration space.
Now $Q_{t}$ comprises the exact positions of all particles in the universe, so observers inside a Bohmian universe cannot possibly measure $Q_{t}$; at best, they can measure a very coarse-grained version of $Q_{t}$. But the point is that even complete information about $Q_{t}$ would not help with distinguishing between typical $\Psi_{t_{0}}$ and $\rho_{0}$ (or between two typical wave functions); and a fortiori, a coarse-grained version does not help.
The essence of the difficulty with distinguishing is an instance of the general problem of deciding between two probability distributions $f_{1}$ and $f_{2}$ (say, on $\mathbb{R}^{n}$) after we are given just one point $X$ that was randomly chosen with distribution either $f_{1}$ or $f_{2}$. If $f_{1}$ and $f_{2}$ are disjoint, then this is possible; if they are nearly disjoint, then it is still possible quite reliably (for example, at the confidence level of $95\%$ often used in statistics). But if $f_{1}$ and $f_{2}$ overlap significantly, and if the observed $X$ lies in the overlap region, then this value of $X$ could have arisen from either one, so we cannot decide which $f$ was used. A reliable decision is possible only if $f_{1}(X)\approx 0$ and $f_{2}(X)$ is significantly non-zero or vice versa. Now in our case, $f_{i}=|\Psi_{i}|^{2}$ and $f_{1}\approx f_{2}$ by (5), so they overlap almost exactly and everywhere; values of $X=Q_{t}$ with $f_{1}(X)\approx 0$ will also have $f_{2}(X)\approx 0$ and therefore, with overwhelming probability, are not going to occur for either $f_{1}$ or $f_{2}$. So, the two cannot be distinguished at any time $t$ (not reliably and in fact, as we will discuss in §3.4, not even unreliably).
This conclusion covers also the possibility that observers might make observations (say, with telescopes) or experiments, as the outcomes of these observations or experiments will be recorded in the configuration of some particles (e.g., when scientists publish their findings). It also covers the possibility of making observations or experiments at several times $t_{1},t_{2},\ldots,t_{k}$, since the outcomes will be recorded and thus can be read off from $Q_{t}$ at $t\geq t_{k}$.
On the other hand, the precise microscopic trajectories of all Bohmian particles may contain more information about $\Psi_{t_{0}}$, but it is known (Dürr et al. 1992) that observers inside a Bohmian universe cannot measure positions at several times without introducing decoherence and thereby changing the trajectory. Therefore, they do not have access to the full trajectories, and the observational indistinguishability just derived remains valid. In other words, there is a known limitation to knowledge about full trajectories—the absolute uncertainty in Bohmian mechanics.
(3) OI in GRWf. We consider GRWf, the GRW theory with a flash ontology in spacetime (e.g., Tumulka 2022b, sec. 3.3.4). A key fact about OI in GRWf is this:
the joint probability distribution of all flashes in the universe up to time $t$ is nearly independent of $\Psi_{t_{0}}$ for typical $\Psi_{t_{0}}\in\mathbb{S}(\mathscr{H}_{0})$. (8)
This follows from Theorem 1 by taking the POVM $E$ to be the POVM governing the joint distribution of all flashes (Tumulka 2022b, sec. 5.1.1 and 7.8). As in BM, observers in a GRWf universe do not have access to the full pattern of flashes, only to a coarse-grained version of it. But even the exact pattern would not provide the information needed for distinguishing a typical $\Psi_{t_{0}}$ from $\rho_{0}$ or from another typical initial wave function, for the same reasons based on Theorem 1 as discussed for BM above.
A difference to the Bohmian case is that the exact history of the universal configuration in BM may provide sufficient information to determine $\Psi_{t_{0}}$, whereas in GRWf even the exact history of all flashes is of no help. Thus, GRWf provides, in a sense, a stronger kind of limitation to knowledge than BM.
Another difference to BM is that inhabitants of a GRWf universe living at time $t$ may very well find out a lot about $\Psi_{t}$. After all, $\Psi_{t}$ is not unitarily evolved from $\Psi_{t_{0}}$ but has collapsed; for example, if we find Schrödinger’s cat alive then its wave function will have collapsed to (approximately) the wave function of a live cat. (Note that this still does not give us new information about $\Psi_{t_{0}}$.)
3.3 Remarks
Let us return to the general result of observation typicality and discuss the order of quantifiers. It has the form “for any $E$ and for most $\Psi$, $s(\Psi,E)$.” This is weaker than the statement “for most $\Psi$ and for any $E$, $s(\Psi,E)$,” which is false in this context. It is easy to confuse the two. The false but stronger statement corresponds to the following:
Super-Strong Observation Typicality
For most quantum states in a high-dimensional Hilbert subspace $\mathscr{H}_{0}$, for any observation $E_{z}$, they cannot be distinguished by $E_{z}$.
Super-strong observation typicality implies observation typicality; it is false because if $E_{z}=|\psi\rangle\langle\psi|$ is a 1d projection in the direction of a particular $\psi\in\mathbb{S}(\mathscr{H}_{0})$, then $\mathbb{P}_{\Psi=\psi}(Z=z)=1$, whereas $\mathbb{P}_{\rho_{0}}(Z=z)=1/d_{0}\approx 0$. The super-strong property requires a uniformity in the class of typical quantum states, while observation typicality is compatible with the possibility that the typicality class depends on the choice of $E_{z}$. However, the weaker statement—observation typicality—is sufficient for our purposes. OI of typical quantum states is true for any chosen experiment, so it is true for the actual experiment that we happen to perform at an arbitrary time.
Another remark concerns alternative arguments for OI using other typicality theorems, specifically canonical typicality (Gemmer et al. 2009, Popescu et al. 2006, Goldstein et al. 2006a) and dynamical typicality as stated in (Teufel et al. 2023). We will discuss their connection to distribution typicality and observation typicality elsewhere (Chen and Tumulka 2024). Here we make two brief comments:
  1. 1.
    Suppose we have access to only local observables pertaining to a small subsystem $\mathscr{S}$ of the universe (less than half the number of degrees of freedom in the universe). Then our observations cannot distinguish typical quantum states $\Psi_{t_{0}}$ of the universe from each other. After all, canonical typicality, applied to a subspace $\mathscr{H}_{R}$ now not given by the micro-canonical energy shell (as in most discussions of canonical typicality) but by $\mathscr{H}_{R}=U_{t}\mathscr{H}_{0}$, asserts that most $\Psi_{t}\in\mathscr{H}_{R}$ have nearly the same reduced density matrix $\rho_{\mathscr{S}}=\operatorname{tr}_{\mathscr{S}^{c}}|\Psi_{t}\rangle\langle\Psi_{t}|$ in $\mathscr{S}$ (where $\mathscr{S}^{c}$ denotes the complement of $\mathscr{S}$). But the Born distribution of any observable in $\mathscr{S}$ depends on $\Psi_{t}$ only through $\rho_{\mathscr{S}}$; thus, the Born distribution accessible to observers like us is nearly independent of $\Psi_{t_{0}}$.
    In this context, the super-strong form of OI is actually true: for a typical quantum state $\Psi_{t_{0}}$, there is no local observation at all that can distinguish it from $\rho_{0}$. After all, once $\operatorname{tr}_{\mathscr{S}^{c}}|\Psi_{t}\rangle\langle\Psi_{t}|\approx\operatorname{tr}_{\mathscr{S}^{c}}(U_{t}\rho_{0}U_{t}^{\dagger})$, all observables in $\mathscr{S}$ have nearly the same Born distribution in $\Psi_{t}$ as in $U_{t}\rho_{0}U_{t}^{\dagger}$.
  2. 2.
    Dynamical typicality means here the statement that most $\Psi_{t_{0}}\in\mathbb{S}(\mathscr{H}_{0})$ have for every $t$ nearly the same Born distribution for all macroscopic observables (and nearly the same as $\rho_{0}$). If any scientist could distinguish $\Psi_{t_{0}}$ from $\rho_{0}$, then she or he could publish the result, say at time $t$, so there would be a macroscopic difference between $\Psi_{t_{0}}$ and $\rho_{0}$, which is impossible by dynamical typicality.
3.4 A Bayesian Analysis
So far, we have argued that no observations can reliably distinguish typical universal quantum states from each other. What about unreliable tests that may nonetheless yield probabilistic information about the quantum state? We can use the previous results to give a Bayesian analysis. Here we provide two quantitative bounds of the probabilistic information obtainable from each observation. It turns out that insofar as the observation is not too improbable, it will not give us substantial probabilistic information about the universal quantum state. Hence, there is a sense that even unreliably, universal quantum states cannot be distinguished.
Suppose the prior probability distribution over quantum states in $\mathscr{H}_{0}$, $\mathbb{P}(\Psi_{t_{0}})$, is given by the Past Hypothesis, i.e., by $u_{0}$. We can show that, if the observation is not too improbable, the posterior probability distribution is very close to the prior.
Corollary 1.
Suppose $d_{0}>1/\varepsilon^{5}$ and $\operatorname{tr}(\rho_{0}\,E_{z})>\varepsilon$. Consider a prior distribution given by $u_{0}$, and let $f(\psi)$ be the density relative to $u_{0}$ of the posterior distribution obtained by Bayesian updating, given that $Z=z$. Then
$1-\varepsilon<f(\psi)<1+\varepsilon$ (9)
for $(1-\varepsilon)$-most $\psi\in\mathbb{S}(\mathscr{H}_{0})$.
This follows from Theorem 1 via a short calculation. [^5] Proof: From $\mathbb{P}(\Psi\in d\psi)=u_{0}(d\psi)$ for every infinitesimal set $d\psi\subset\mathbb{S}(\mathscr{H}_{0})$ and $\mathbb{P}(Z=z|\Psi)=\langle\Psi|E_{z}|\Psi\rangle$, we find that $\mathbb{P}(\Psi\in d\psi,Z=z)=u_{0}(d\psi)\langle\psi|E_{z}|\psi\rangle$ and $\mathbb{P}(Z=z)=\int_{\mathbb{S}(\mathscr{H}_{0})}u_{0}(d\phi)\langle\phi|E_{z}|\phi\rangle=\operatorname{tr}(\rho_{0}E_{z})$. Therefore, $u_{0}(d\psi)\,f(\psi)=\mathbb{P}(\Psi\in d\psi|Z=z)=u_{0}(d\psi)\langle\psi|E_{z}|\psi\rangle/\operatorname{tr}(\rho_{0}E_{z})$. By Theorem 1 for $(1-\varepsilon)$-most $\psi$, $\bigl{|}\langle\psi|E_{z}|\psi\rangle-\operatorname{tr}(\rho_{0}E_{z})\bigr{|}<\varepsilon^{2}$, and thus $\bigl{|}\langle\psi|E_{z}|\psi\rangle/\operatorname{tr}(\rho_{0}E_{z})-1\bigr{|}<\varepsilon^{2}/\operatorname{tr}(\rho_{0}E_{z})<\varepsilon$. We can also reformulate the result in terms of the probability of any set $S\subseteq\mathbb{S}(\mathscr{H}_{0})$ instead of the density function $f$:
Corollary 2.
Suppose $d_{0}>1/\varepsilon^{5}$ and $\operatorname{tr}(\rho_{0}\,E_{z})>\varepsilon$. For any subset $S\subseteq\mathbb{S}(\mathscr{H}_{0})$:
$\mathbb{P}(\Psi_{t_{0}}\in S\,|\,Z=z)\in\Bigl{[}\mathbb{P}(\Psi_{t_{0}}\in S)-2\varepsilon,\mathbb{P}(\Psi_{t_{0}}\in S)+3\varepsilon\Bigr{]}$ (10)
This follows from Corollary 1 via standard arguments of probability and integration theory. [^6] Proof: Let $M\subseteq\mathbb{S}(\mathscr{H}_{0})$ be the set where (9) holds. By Corollary 1, $u_{0}(M)\geq 1-\varepsilon$. Thus, writing $M^{c}=\mathbb{S}(\mathscr{H}_{0})\setminus M$ for the complement of $M$ and using (9), $\mathbb{P}(\Psi\in S|Z=z)=\int_{S}u_{0}(d\psi)\,f(\psi)\geq\int_{S\cap M}u_{0}(d\psi)\,f(\psi)\geq\int_{S\cap M}u_{0}(d\psi)\,(1-\varepsilon)=(1-\varepsilon)\,u_{0}(S\cap M)\geq(1-\varepsilon)\,(u_{0}(S)-u_{0}(M^{c}))\geq(1-\varepsilon)\,(u_{0}(S)-\varepsilon)\geq u_{0}(S)-2\varepsilon$. On the other hand, since $f$ is normalized, $\int_{\mathbb{S}(\mathscr{H}_{0})}u_{0}(d\psi)\,f(\psi)=1$, we have that $\int_{M^{c}}u_{0}(d\psi)\,f(\psi)=1-\int_{M}u_{0}(d\psi)\,f(\psi)\leq 1-\int_{M}u_{0}(d\psi)\,(1-\varepsilon)=1-(1-\varepsilon)\,u_{0}(M)\leq 1-(1-\varepsilon)^{2}<2\varepsilon$. Thus, using (9) again, $\mathbb{P}(\Psi\in S|Z=z)=\int_{S}u_{0}(d\psi)\,f(\psi)=\int_{S\cap M}u_{0}(d\psi)\,f(\psi)+\int_{S\cap M^{c}}u_{0}(d\psi)\,f(\psi)\leq\int_{S\cap M}u_{0}(d\psi)\,(1+\varepsilon)+\int_{M^{c}}u_{0}(d\psi)\,f(\psi)<(1+\varepsilon)\,u_{0}(S\cap M)+2\varepsilon\leq(1+\varepsilon)u_{0}(S)+2\varepsilon\leq u_{0}(S)+3\varepsilon$.
The assumption $d_{0}>\varepsilon^{-5}$ is reasonable; even when $\varepsilon=10^{-200}$, the condition is still satisfied as $\varepsilon^{-5}=10^{1000}\leq 10^{10^{20}}\leq d_{0}$. In this case, the posterior probability distribution differs minimally from the prior and is almost independent of any observation. This conclusion requires the assumption that the observation is not too improbable.
Hence, distribution typicality entails not only OI, but also that, in general, observations reveal very little about the initial quantum state. Notice how this fact fails in classical mechanics. Observations, even when they are probable, can reveal a great deal about the classical initial state. For example, observing the Tails outcome of a classical coin-flip experiment can rule out half the initial conditions, namely those that predict a Heads outcome. Observations cannot reveal so much in quantum mechanics because of distribution typicality and observation typicality, two facts whose analogues fail in classical statistical mechanics.
4 Comparisons with Known Results
In this section, we briefly compare our results for OI with some known results in philosophy of science and foundations of physics.
In quantum mechanics, there are several results about limitation to knowledge about the quantum state. For a survey, see (Tumulka 2022b, ch.5). Here we highlight a few:
  • Because of the Main Theorem about POVMs, we cannot measure the wave function of a given system. Relatedly, because of the no-cloning theorem, we cannot reliably copy the wave function of a given system.
  • It can be shown (Tumulka 2022a) that in any ontological theory of quantum mechanics, it is impossible to measure the ontic state. Thus, whatever the true theory of quantum mechanics may be, there must be facts that cannot be empirically determined.
  • Because of the Pusey-Barrett-Rudolph (PBR) theorem (Pusey et al. 2012), two different ensembles of wave functions with the same density matrix are physically distinct. However, they are observationally indistinguishable by all possible observations directly on the ensembles.
The last item also asserts a kind of OI, but differs from our result here in that it provides a condition (equal density matrix) under which two situations (ensembles of wave functions) are indistinguishable, whereas we show here that most wave functions from $\mathscr{H}_{0}$ are indistinguishable. Moreover, all three items are compatible with substantive learning about the actual quantum state. One may still hope to rule out a significant fraction of quantum states based on some observation. Moreover, a Bayesian can, in general, update their probability distribution based on observations after which their posterior may be highly peaked. In contrast, observation typicality tells us that observations in a typical quantum universe do not rule out more than a tiny fraction of universal quantum states, and (if the observation is not too improbable) do not have a substantial influence on a uniform prior probability distribution about initial quantum states. It stays nearly uniform.
We now compare our results to a much discussed underdetermination result in the philosophy of general relativity (GR). Manchak (2009) showed that for almost every spacetime model of GR, there exists another spacetime that is physically distinct (in the sense of having a different global structure) but observationally indistinguishable from it. [^7] For some philosophical discussions about this result, see Beisbart (2009), Norton (2011), Butterfield (2012), and Cinti and Fano (2021).
Both Manchak’s result and ours are compatible with a number of inductive hypotheses. In Manchak’s case, one can assume that the same law of nature (Einstein’s equation of GR) applies to all spacetime points. Moreover, Manchak’s result is compatible with any set of local conditions (conserved under local isometries). In our case, one can also assume that the same laws of nature (within quantum theory) apply to the entire model, so that we are considering the same Hilbert space $\mathscr{H}$, the same Hamiltonian $H$, and the same subspace $\mathscr{H}_{0}$.
There are important differences too. First, Manchak’s argument shows the existence of an OI counterpart for almost every model of GR, while our argument shows that almost every model is an OI counterpart to almost every other model of QM. Hence, his result is compatible with the fact that we can observationally rule out a significant fraction of possible universes while ours is not. Second, Manchak’s proof uses a “cut-and-paste” method to construct special models of GR that are OI from the original model, while ours uses the probabilistic method to show that OI must hold among typical models of QM with generic features.
Related to the second point, Manchak’s argument has been challenged on philosophical grounds that the constructed spacetimes are not “physically reasonable,” because some people regard the “cut-and-paste” models as implausible candidates for the actual universe. (For example, see Cinti and Fano (2021).) One may respond, as Manchak does, that “the theorem can have physical relevance even if the particular model constructed in the proof does not” (Manchak 2011, p.7). We do not take a stance on that issue, but merely point out that the objection of “physical unreasonableness” does not even work as a prima facie objection against our argument for OI in QM. We do not show that almost any model has at least one specially constructed counterpart. Rather, we show that typical (generic) models of QM in a high-dimensional Hilbert space are OI from each other.
It is interesting to note that GR and QM, two pillars of modern physics, both lead to OI.
5 Philosophical Implications
We highlight some potential implications of observation typicality for philosophy of science. First, observation typicality places severe limitations to knowledge in a quantum universe. In a large universe, no experiment can distinguish the overwhelming majority of universal quantum states. This is an in-principle limitation that cannot be overcome by new technologies.
Second, the flip side is that, concerning the observable properties of the universal quantum state, we may know a lot about them because of the typicality results. For any observation, almost all universal quantum states look the same as the normalized projection $\rho_{0}$. Assuming our universe is typical among those that satisfy the PH, even though we do not know which it is exactly, we know there is uniformity in their predictions—they are very nearly the same as the probabilistic prediction of the Wentaculus density matrix. Fixing the PH nearly fixes the probabilistic predictions of typical quantum states.
It provides a new perspective on the predictive power of initial condition laws such as the PH. When the initial condition law delivers a high-dimensional subspace in the full Hilbert space, the precise choice of a typical quantum state makes almost no difference to the probabilistic prediction. For any observation, typical individual quantum states predict with nearly the same probability distribution. This raises an interesting question, which we leave to future work, whether typical choices of the initial condition, in some sense to be defined (e.g. typical subspace), lead to nearly the same probability distributions.
Third, observation typicality may conflict with positivism, the idea that a statement is unscientific or meaningless if it cannot be tested experimentally, and a variable is not well-defined if it cannot be measured. Given the PBR theorem, and the various solutions to the quantum measurement problem, we have good reasons to accept that the universal quantum state is an objective feature of the universe. Due to observation typicality, however, it is undetertermined by observational evidence. This is ironic, to echo a point made by Cowan and Tumulka (2016): positivism is sometimes supported by examples from quantum mechanics, and yet quantum mechanics tells us that the central object in a quantum universe is well-defined, objective, but non-measurable, which undermines positivism.
Finally, our result has potential implications for realism about the quantum state. On realism, it is an open question what the universal quantum state represents and how best to understand its objectivity and reality. (For a survey, see Chen (2019).) Within realism, there are two main approaches. The first is to regard it as a physical thing, like a physical field, that exists as part of the basic building blocks of the universe (Albert 1996, Ney 2021). The second is to regard it as a physical law, like the Hamiltonian function in classical mechanics, that tells physical things (such as particles and other fields) how to move (Goldstein and Zanghì 2013).
If the quantum state is a physical thing, OI suggests an in-principle limit to how precisely it can be measured: we cannot gain substantive information about which quantum state is the actual one. Physical objects in classical mechanics do not have this character. They can be measured to arbitrary precision. Almost all quantum states of the universe will appear the same to us. This should be challenging to the empiricists who hold out hope that the basic building blocks of the universe should be in-principle measurable.
Perhaps the result is less surprising if the universal quantum state has the status of a physical law. Since it is underdetermined by observational evidence, we have to use theoretical virtues, such as simplicity and elegance, to pin it down. The theoretical flavor of how we choose the universal quantum state is similar to how we choose other physical laws, whose precise forms are constrained but still vastly underdetermined by observational evidence.
Realism about the universal quantum state also leaves open whether it is necessarily a pure state represented by a wave function, or possibly an impure (“mixed”) one represented by a density matrix. Following Chen (2021, 2024b), we call the first option wave function realism (WFR) and the second density matrix realism (DMR). It is already known that WFR and DMR are empirically equivalent, given appropriate choices of the universal quantum states. Observation typicality suggests that it is even harder to distinguish the two theories. Not only are typical individual wave functions OI from each other, but they are also OI from the normalized projection, an impure density matrix. One may regard it as a potential reason to prefer DMR over WFR. If typical pure states $\Psi_{t_{0}}$ compatible with the Past Hypothesis do not lead to different predictions, why not just use the Wentaculus density matrix $\rho_{0}$ instead? After all, $\rho_{0}$ may be regarded as much simpler than any typical $\Psi_{t_{0}}$ even though they are nearly predictively equivalent.
Here we do not take a firm stance on these implications, merely exploring how they may be suggested by our results.
6 Conclusion
We have shown that the central object in quantum theory—the quantum state of the universe—is most likely hidden from observation, because typical quantum states in a high-dimensional Hilbert subspace $\mathscr{H}_{0}$ are observationally indistinguishable from the density matrix $\rho_{0}=P_{0}/d_{0}$ and thus from each other. In fact, no observation will yield any substantial information at all, even partial or probabilistic, about $\Psi_{t_{0}}$. That is in a sense the strongest known result about a limitation to knowledge in a quantum universe. Assuming the universe we inhabit is typical, our observational data alone will tell us very little about exactly which one it is that we inhabit. Nature is more secretive than we have realized.
Acknowledgment. We thank Sheldon Goldstein for extensive feedback and for suggesting the term “observation typicality.” For helpful discussions, we also thank Emily Adlam, David Albert, Fabio Anza, Patricio Avila Cardenas, Jeff Barrett, Jeremy Butterfield, Craig Callender, Branden Fitelson, Hans Halvorson, Alan Hájek, Tyler Hildebrand, Christopher Jarzynski, Joel Lebowitz, JB Manchak, Kerry McKenzie, Marcos Rigol, Daniel Rubio, Tony Short, Stefan Teufel, Isaac Wilhelm, Nicole Yunger Halpern, the UC San Diego Philosophy of Physics Group, and the participants at the 2024 Quantum Thermodynamics Workshop at the University of Maryland, College Park.
Funding. EKC is supported by Grant 63209 from the John Templeton Foundation. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation.
Conflict of interests. The authors declare no conflict of interest.

### v2

@@
 Typical Quantum States of the Universe are Observationally Indistinguishable
-Abstract: This paper is about the epistemology of quantum theory.
-We establish a new result about a limitation to knowledge of its central object—the quantum state of the universe.
-We show that, if the universal quantum state can be assumed to be a typical unit vector from a high-dimensional subspace of Hilbert space (such as the subspace defined by a low-entropy macro-state as prescribed by the Past Hypothesis), then no observation can determine (or even just narrow down significantly) which vector it is.
-Typical state vectors, in other words, are observationally indistinguishable from each other.
-Our argument is based on a typicality theorem from quantum statistical mechanics.
-We also discuss how theoretical considerations that go beyond the empirical evidence might bear on this fact and on our knowledge of the universal quantum state.
-Key words: limitation to knowledge; empirical equivalence; past hypothesis; quantum statistical mechanics.
+Abstract: We establish three new impossibility results regarding our knowledge of the quantum state of the universe—a central object in quantum theory.
+We show that, if the universal quantum state is a typical unit vector from a high-dimensional subspace $\mathscr{H}_{0}$ of Hilbert space $\mathscr{H}$ (such as the one defined by a low-entropy macro-state as prescribed by the Past Hypothesis), then no observation can determine or just significantly narrow down which vector it is.
+In other words, the overwhelming majority of possible state vectors are observationally indistinguishable from each other (and from the density matrix of $\mathscr{H}_{0}$).
+Moreover, we show that for any observation that isn’t too unlikely and most pairs of unit vectors from $\mathscr{H}_{0}$, the observation will not significantly favor one vector over the other.
+We further show that the uniform distribution over the unit sphere in $\mathscr{H}_{0}$, after Bayesian updating in the light of any observation that isn’t too unlikely, is still extremely close to uniform.
+Our arguments rely on a typicality theorem from quantum statistical mechanics.
+We also discuss how theoretical considerations beyond empirical evidence might inform our understanding of this fact and our knowledge of the universal quantum state.
+Key words: limitation to knowledge; empirical equivalence; past hypothesis; quantum statistical mechanics; quantum measurement; measure concentration.
 1 Introduction
-This paper is about the epistemology of quantum theory.
-We establish a new result about a limitation to knowledge of the quantum state of the universe.
-We show that typical universal quantum states are observationally indistinguishable, in the sense that no observation can distinguish typical universal quantum states in a high-dimensional subspace of the Hilbert space, such as the subspace characterized by a low-entropy macro-state as prescribed by the Past Hypothesis (see §2.2 and (Albert 2000, Goldstein et al.
-2020)).
+We establish three new impossibility results regarding our knowledge of the quantum state of the universe—a central object in quantum theory.
+We show that typical universal quantum states are observationally indistinguishable, in the sense that no observation can distinguish typical universal quantum states in a high-dimensional subspace of the Hilbert space, such as the one prescribed by the Past Hypothesis (see §2.2 and (Albert 2000, Goldstein et al.
+2020)), unless we somehow already knew the quantum state of the universe.
 We call this fact observation typicality.
 Observation typicality follows from a result in quantum statistical mechanics called distribution typicality (Reimann 2007, Teufel et al.
-2023): for any observable, most quantum states in a high-dimensional Hilbert subspace lead to very nearly the same probability distribution for that observable.
-We show that this implies observation typicality.
-Observation typicality is a surprisingly strong result.
+2023): for any observable, most quantum states in a high-dimensional Hilbert subspace $\mathscr{H}_{0}$ lead to very nearly identical probability distributions for that observable.
+Roughly speaking, that is, most wave functions from a high-dimensional $\mathscr{H}_{0}$ look the same.
+We assume here that there is a wave function $\Psi_{t}$ of the universe and that at the initial time $t_{0}$ of the universe (say, at the big bang), $\Psi_{t_{0}}$ had to lie in a particular subspace $\mathscr{H}_{0}$ of the Hilbert space $\mathscr{H}$ of the universe (say, corresponding to a low-entropy macro-state, §2.2).
+Observation typicality then asserts that our empirical data at any time $t$ will reveal very little about $\Psi_{t_{0}}$ (beyond its membership in $\mathscr{H}_{0}$, which we take as known), even if we assume that any observable (including any POVM) can be measured.
+Observation typicality is a surprisingly strong result, yet it has gone unnoticed so far.
 There are other known examples of epistemic limitations in physics.
@@
 (2004) already proved that it is impossible to measure a wave function precisely.
-Observation typicality is a further step in this direction, and it delivers a stronger conclusion—even an imprecise measurement is impossible.
-No observation, as long as it does not disconfirm the physical theory, will give us any substantive information about the quantum state.
-In this sense, limitation to knowledge is more pervasive and persistent in a quantum universe than has been recognized.
-A statement about most quantum states in the full Hilbert space $\mathscr{H}$ of the universe has limited applicability because it may fail for the quantum states of interest.
+Observation typicality takes this limitation a step further, delivering an even stronger conclusion: typically, even an imprecise measurement is impossible.
+No observation—provided it does not disconfirm the physical theory—can yield substantive information about the quantum state.
+In this sense, the limitation to knowledge is more pervasive and persistent in a quantum universe than previously recognized.
+We note that a statement about most quantum states in the full Hilbert space $\mathscr{H}$ of the universe would have limited applicability because it may fail for the quantum states of interest.
 For example, most quantum states in $\mathscr{H}$ are in thermal equilibrium, but the actual quantum state of our universe is not.
@@
 We assume that $\mathscr{H}_{0}$ has finite but huge dimension $d_{0}$ (while $\mathscr{H}$ may have finite or infinite dimension).
-Observation typicality then holds for most quantum states $\Psi$ in $\mathscr{H}_{0}$.
-We can use the result while accepting certain inductive hypotheses about the physical laws of the universe, including those pertaining to initial conditions, according to which our universe is a typical member of the possible universes compatible with such laws.
-We assume here that there is a wave function $\Psi_{t}$ of the universe and that at the initial time $t_{0}$ of the universe (say, at the big bang), $\Psi_{t_{0}}$ had to lie in a particular subspace $\mathscr{H}_{0}$ of the Hilbert space $\mathscr{H}$ of the universe (say, corresponding to a low-entropy macro-state, as prescribed by the Past Hypothesis).
-Observation typicality asserts that our empirical data at any time $t$ will reveal very little about $\Psi_{t_{0}}$ (beyond its membership in $\mathscr{H}_{0}$, which we take as known) because most $\Psi_{t_{0}}$ from $\mathscr{H}_{0}$ lead to very nearly the same probability distribution for these empirical data.
-In other words, typical initial quantum states of the universe are observationally indistinguishable from each other.
-If the time evolution is unitary, then the same arguments yield that also $\Psi_{t}$ cannot be distinguished empirically from typical alternative vectors in the appropriate subspace, i.e., in $U_{t}\mathscr{H}_{0}$ with $U_{t}=\exp(-iH(t-t_{0}))$ and $H$ the Hamiltonian of the universe.
-The analogous statement fails in classical mechanics.
-Observation typicality holds in quantum mechanics because, roughly speaking, quantum states can be in superpositions, and the superposition weights of any particular observation given by typical quantum states are nearly the same.
-The latter is related to the mathematical fact, known as the concentration of measure phenomenon (the situation where random quantities become nearly deterministic), which applies in any high-dimensional Hilbert space.
-We present two main results.
-First (§2 and §3), that a typical unit vector $\Psi_{t_{0}}\in\mathscr{H}_{0}$ cannot be reliably distinguished from the density matrix $\rho_{0}$ associated with a uniform distribution in $\mathscr{H}_{0}$, i.e., $\rho_{0}=P_{0}/d_{0}$ with $P_{0}$ the projection operator to $\mathscr{H}_{0}$ and $d_{0}=\dim\mathscr{H}_{0}$; as a consequence, typical unit vectors cannot be reliably distinguished from each other.
-The second result (§3.4) is even stronger and asserts that even unreliably, they cannot be distinguished; that is, our empirical data do not even yield (any significant amount of) partial or probabilistic information about $\Psi_{t_{0}}$.
-This result can be expressed in terms of Bayesian credences as follows: if we start from the uniform probability distribution $u_{0}$ over the unit sphere $\mathbb{S}(\mathscr{H}_{0})$ (containing all normalized wave functions) in $\mathscr{H}_{0}$ as the prior distribution, and if we update our credences in a Bayesian way on the basis of our empirical observation (that is not too improbable), then the updated distribution will still be very nearly uniform over $\mathbb{S}(\mathscr{H}_{0})$.
+[^3] If our universe initially had a finite volume and occupied a finite energy interval $[E_{1},E_{2}]$, it is confined to a Hilbert space $\mathscr{H}_{[E_{1},E_{2}]}$ of finite (albeit realistically very large) dimension.
+See §2.2.
+We can use the result while accepting certain physical laws, including those pertaining to initial conditions of the universe, according to which our universe is a typical member of the possible universes compatible with such laws.
+It is well known that if we pick two vectors $\Psi_{A},\Psi_{B}$ purely randomly from the unit sphere of a high-dimensional $\mathscr{H}_{0}$, then with overwhelming probability they will be nearly orthogonal, and thus there exists a quantum observable $P$ that distinguishes them reliably.
+However, this fact does not affect the question we are studying here because the choice of $P$ depends on (at least one of) $\Psi_{A},\Psi_{B}$, while we do not assume we already know the wave function of the universe.
+The situation that we do not allow the observable to depend on the quantum state is reflected in formal statements by the order of quantifiers such as some, most, and every.
+To clarify the order of quantifiers in our result: By observation typicality, we mean that for every observation, most quantum states are not distinguishable by that observation.
+The alternative order—where we first quantify over most states and then over every observation—would imply an even stronger claim, which we call super-strong observation typicality.
+The latter claim is generally false, though it holds for local observations because of the more well-known result of canonical typicality, as we discuss in Chen and Tumulka (2025).
+In classical mechanics, the statement analogous to observation typicality fails.
+It holds in quantum mechanics because, roughly speaking, quantum states can be in superpositions, and the superposition weights of any particular observation given by typical quantum states are nearly identical.
+The latter is a version of the mathematical fact known as the concentration-of-measure phenomenon (the situation where random quantities become nearly deterministic) and occurs in high-dimensional Hilbert spaces.
+We note further that the same arguments used to establish observation typicality yield that, if the time evolution is unitary, also $\Psi_{t}$ (instead of $\Psi_{t_{0}}$) cannot be distinguished empirically from typical alternative vectors in the appropriate subspace, i.e., in $U_{t}\mathscr{H}_{0}$ with $U_{t}=\exp(-iH(t-t_{0}))$ and $H$ the Hamiltonian of the universe.
+Our conclusions are based on three mathematical results.
+First (§3.2.1), that a typical unit vector $\Psi_{t_{0}}\in\mathscr{H}_{0}$ cannot be reliably distinguished from the density matrix $\rho_{0}$ associated with a uniform distribution in $\mathscr{H}_{0}$, i.e., $\rho_{0}=P_{0}/d_{0}$ with $P_{0}$ the projection operator to $\mathscr{H}_{0}$ and $d_{0}=\dim\mathscr{H}_{0}$.
+Consequently, typical unit vectors cannot be reliably distinguished from each other.
+By contrast, the second result (§3.2.2) is in a sense stronger but does require that the observation in question be not too improbable.
+It asserts that even unreliably, the states cannot be distinguished, which conveys that such empirical data do not even yield (any significant amount of) partial or probabilistic information about $\Psi_{t_{0}}$.
+The third result (§3.2.3) is expressed in Bayesian terms: if we start from the uniform prior $u_{0}$ over the unit sphere $\mathbb{S}(\mathscr{H}_{0})$ (containing all normalized wave functions) in $\mathscr{H}_{0}$ and update in a Bayesian way upon an observation that is not too improbable, then the posterior distribution will still be very nearly uniform over $\mathbb{S}(\mathscr{H}_{0})$.
 Some empirical observations might, of course, change our mind about which subspace $\mathscr{H}_{0}$ is the correct one.
-Since we regard here the specification of $\mathscr{H}_{0}$ as part of the given theory, such an observation would amount to a disconfirmation of the theory.
+Since we regard here the specification of $\mathscr{H}_{0}$ (in the Past Hypothesis) as part of the given theory, such an observation would amount to a disconfirmation of the theory.
 This leads to the interesting question of which kind of observations should be regarded as disconfirming the theory.
-We will look more into this question elsewhere (Chen and Tumulka 2024), take here for granted that the theory is correct and not getting disconfirmed, and remark only briefly that since the specification of $\mathscr{H}_{0}$ has the status of a law of nature, we may expect it to be simple and natural in order to be convincing; as a consequence, we would not be inclined to adjust the choice of $\mathscr{H}_{0}$ in complicated ways just to adapt it to some random-looking features of our world (such as, e.g., the shape of Ireland).
-Our goal is to further the discussion about the epistemology of physical theories and highlight the limitations of (direct) empirical knowledge in quantum universes.
-Insofar as we have detailed knowledge about the universal quantum state, it is even more theoretical than has been recognized.
-Our result should be of interest to anyone interested in the epistemological implications of physics.
-We highlight several distinctive features of the present approach:
-  • Observation typicality applies to all interpretations of quantum mechanics.
-It is also compatible with generous inductive assumptions about what we can infer based on available evidence.
-For example, it is compatible with our knowledge of the dynamical laws of quantum mechanics, the Past Hypothesis, and the aggregation of individual observations into a collection.
-  • Our argument relies on the concentration of measure phenomenon instead of special constructive techniques, such as the cut-and-paste method used in Manchak (2009)’s argument for observational indistinguishability in general relativity.
-The standard “physical unreasonableness” objections against arguments based on special constructions do not apply to our argument, which establishes observational indistinguishability for generic pairs of models.
-  • Our claims are based on quantitative bounds (recently proven mathematically) on how large the deviations of certain probabilities can at most become for which fraction of all wave functions in a high-dimensional Hilbert subspace.
+We will look more into this question elsewhere (Chen and Tumulka 2025), take here for granted that the theory is correct and not getting disconfirmed, and remark only briefly that if the specification of $\mathscr{H}_{0}$ has the status of a law of nature, we may expect it to be simple and natural in order to be convincing (Loewer 2024, Chen 2024c); as a consequence, we would not be inclined to adjust the choice of $\mathscr{H}_{0}$ in complicated ways just to adapt it to some random-looking features of our world (such as, e.g., the shape of Ireland).
+Our results show that insofar as we have detailed knowledge about the universal quantum state, it is even more theoretical than has been recognized.
+They should be of interest to anyone interested in the epistemological implications of physics.
+This paper is organized as follows.
 In §2, we formulate distribution typicality and explain how it applies to the special case of the Past Hypothesis.
-In §3, we clarify our criterion for observational indistinguishability, derive observation typicality from distribution typicality, apply the fact to the universal quantum state, and provide a Bayesian analysis.
+In §3, we clarify our criterion for observational indistinguishability, derive observation typicality from distribution typicality, and provide a Bayesian analysis.
 In §4, we compare observation typicality to known results about epistemic limitations in physics.
 In §5, we discuss potential philosophical implications.
-For more detail, see the longer version, Chen and Tumulka (2024).
+For more detail, including the application of observation typicality to subsystem quantum states of the universe, see Chen and Tumulka (2025).
+In §6, we conclude.
 2 Distribution Typicality
-In the 21st century, many exciting new results have been proven in quantum statistical mechanics and quantum thermodynamics, leading to improved understanding in the thermodynamic behavior of a closed quantum system in a pure state (Gemmer et al.
+In the 21st century, many exciting new results have been proven in quantum statistical mechanics and quantum thermodynamics, leading to improved understanding of the thermodynamic behavior of a closed macroscopic quantum system in a pure state (Gemmer et al.
 2009, Tasaki 2016, Mori et al.
 2018).
-Such results often assume an “individualist” attitude, according to which an individual quantum system in a pure state, as opposed to an ensemble, is studied with respect to its thermodynamic properties.
+Such results often assume an “individualist” attitude, according to which an individual quantum system in a pure state, as opposed to either a density matrix or an ensemble of systems, is studied with respect to its thermodynamic properties.
 These results often concern typical properties of those states.
-[^3] There is a growing philosophical literature about typicality and its role in physical theories.
-See, for example, Goldstein (2012), Reichert (2018), Wilhelm (2022), Lazarovici (2023), Hubert (2024), (Frigg and Werndl 2024, sec.
+[^4] The concept of typicality and its role in physical theories have received growing attention in the philosophical literature.
+For some recent examples, see Goldstein (2012), Lazarovici and Reichert (2015), Allori (2020), Wilhelm (2022), Lazarovici (2023), Hubert (2024), (Frigg and Werndl 2024, sec.
 4.5), and the references therein.
 We shall focus on distribution typicality in this paper.
-In §3, we show that distribution typicality implies observation typicality.
+In §3, we show that distribution typicality implies observation typicality for orthodox quantum mechanics (OQM), Bohmian mechanics (BM), the Ghirardi-Rimini-Weber spontaneous collapse theory (GRW), and Everettian quantum mechanics (EQM) (Tumulka 2022b, Dürr and Teufel 2009, Ghirardi and Bassi 2020, Barrett 2023).
 In this section, we formulate distribution typicality and specialize it to the Past Hypothesis.
 2.1 The Key Theorem
-Before stating the key theorem, we highlight its implications for orthodox quantum mechanics (OQM), Bohmian mechanics (BM), and the Ghirardi-Rimini-Weber spontaneous collapse theory (GRW) (Tumulka 2022b, Dürr and Teufel 2009, Ghirardi and Bassi 2020).
-Distribution typicality entails that, in OQM, if an observer at time $t$ conducts an experiment, the probability distribution of the outcome is nearly independent of $\Psi_{t_{0}}$; in BM, at any time $t$, the probability distribution of the Bohmian configurations $Q_{t}$ of the universe is nearly independent of $\Psi_{t_{0}}$; in GRW with a flash ontology, for any time $t>t_{0}$, the probability distribution of the pattern of flashes up to $t$ is nearly independent of $\Psi_{t_{0}}$.
-We will elucidate in §3 why this near-independence entails the impossibility for observers according to OQM, BM, or GRW to distinguish among different $\Psi_{t_{0}}$.
 Terminology.
 Recall that $\mathbb{S}(\mathscr{H}_{0})$ is the unit sphere in the subspace $\mathscr{H}_{0}$, $d_{0}=\dim\mathscr{H}_{0}$, and $u_{0}$ the uniform probability distribution over $\mathbb{S}(\mathscr{H}_{0})$; that is, for any subset $S$ of $\mathbb{S}(\mathscr{H}_{0})$, $u_{0}(S)$ is the surface area of $S$, normalized by dividing through the surface area of $\mathbb{S}(\mathscr{H}_{0})$.
-We say that a statement $s(\Psi)$ is true for $(1-\varepsilon)$-most $\Psi\in\mathbb{S}(\mathscr{H}_{0})$ if and only if the set $S$ of $\Psi\in\mathbb{S}(\mathscr{H}_{0})$ satisfying $s(\Psi)$ has $u_{0}(S)\geq 1-\varepsilon$ (equivalently, if $S$ has at least the fraction $1-\varepsilon$ of the surface area, or if a purely random point on the sphere has probability $\geq 1-\varepsilon$ to lie in $S$).
+We say that some statement about the wave function $\Psi$, $s(\Psi)$, is true for $(1-\varepsilon)$-most $\Psi\in\mathbb{S}(\mathscr{H}_{0})$ if and only if the set $S$ of $\Psi\in\mathbb{S}(\mathscr{H}_{0})$ satisfying $s(\Psi)$ has $u_{0}(S)\geq 1-\varepsilon$ (equivalently, if $S$ has at least the fraction $1-\varepsilon$ of the surface area, or if a purely random point on the sphere has probability $\geq 1-\varepsilon$ to lie in $S$).
 Averages on $\mathbb{S}(\mathscr{H}_{0})$ will also be taken with respect to $u_{0}$.
@@
 POVM.
-POVM stands for positive-operator-valued measure, which is a generalization of the concept of an observable as given by a self-adjoint operator.
-Technically, a POVM on a Hilbert space $\mathscr{H}$ is a family of positive operators $E_{z}$ on $\mathscr{H}$ that add up to the identity operator.
+A POVM (positive-operator-valued measure) is a generalization of the notion of an observable represented by a self-adjoint operator.
+In the simple case with discrete outcomes, a POVM on a Hilbert space $\mathscr{H}$ is a family of positive operators $E_{z}$ on $\mathscr{H}$ that add up to the identity operator.
+More generally, a POVM is a map $E$ from the measurable subsets of an outcome space $\mathscr{Z}$ to the set of positive operators on $\mathscr{H}$ satisfying normalization, $E(\mathscr{Z})=I$, and countable additivity.
 POVMs have an important representational role in quantum mechanics, given by the Main Theorem about POVMs: for every quantum physical experiment $\mathscr{E}$ on a quantum system $\mathscr{S}$ whose possible outcomes lie in a space $\mathscr{Z}$, there exists a POVM E on $\mathscr{Z}$ such that, whenever $\mathscr{S}$ has wave function $\Psi$ at the beginning of $\mathscr{E}$, the random outcome $Z$ has probability distribution given by
 $\mathbb{P}(Z=z)=\langle\Psi|\,E_{z}\,|\Psi\rangle.$ (1)
-See Dürr et al. (2004) and Tumulka (2022b) for proofs.
+where $E_{z}$, the image of $\{z\}$ under $E$, is an element of the POVM $E$. See Dürr et al. (2004) and Tumulka (2022b) for proofs.
 The Main Theorem about POVMs provides a generalization of the Born rule to arbitrary experiments (instead of ideal quantum measurements).
@@
 Theorem 1 (Distribution Typicality).
-Let $\mathscr{H}_{0}$ be an arbitrary subspace with finite dimension $d_{0}$ in $\mathscr{H}$, let the density matrix $\rho_{0}=P_{0}/d_{0}$ be the normalized projection to $\mathscr{H}_{0}$, and let the operator $E_{z}$ be an element of an arbitrary POVM on $\mathscr{H}$.
+Let $\mathscr{H}_{0}$ be an arbitrary subspace with finite dimension $d_{0}$ in $\mathscr{H}$, let the density matrix $\rho_{0}=P_{0}/d_{0}$ be the normalized projection to $\mathscr{H}_{0}$, and let $B$ be any self-adjoint operator on $\mathscr{H}$.
 Then for every $\varepsilon>0$, for $(1-\varepsilon)$-most $\Psi\in\mathbb{S}(\mathscr{H}_{0})$,
-$\Bigl{|}\langle\Psi|\,E_{z}\,|\Psi\rangle-\operatorname{tr}(\rho_{0}\,E_{z})\Bigr{|}\leq\frac{1}{\sqrt{\varepsilon d_{0}}}\,.$ (2)
-Theorem 1 is a particular case of Theorem 3 proven by Teufel et al. (2023). It can also be derived from a theorem proven by Reimann (2007).
-It can be shown that $\operatorname{tr}(\rho_{0}E_{z})$ is exactly the average of $\langle\Psi|E_{z}|\Psi\rangle$ over $\mathbb{S}(\mathscr{H}_{0})$ (using the uniform distribution $u_{0}$); thus (2) expresses that $\langle\Psi|E_{z}|\Psi\rangle$ is close to its average value, provided that the right-hand side is small, which will be the case whenever
-$d_{0}\gg\frac{1}{\varepsilon}\,.$ (3)
-Inequality (2) is an example of the “concentration of measure phenomenon,” which can be expressed as saying that some random quantity is nearly deterministic (i.e., very probably close to a deterministic value).
-Here, we have that for any observation represented by an element in a POVM, typical quantum states in a high-dimensional Hilbert space have very nearly the same Born probability distribution for that observation.
-Let us clarify and elaborate on this result.
-Example.
-A macroscopic system with $N=10^{20}$ particles is in macro-state $\mu$.
-The dimension of the Hilbert space $\mathscr{H}_{0}$ of the system scales exponentially with $N$.
-Let us suppose $d_{0}=10^{10^{20}}$.
-Consider a very small error $\varepsilon=10^{-200}$.
-Inequality (2) tells us that, for any quantum experiment with observational outcome $z$, for $(1-\varepsilon)$-most pure states in $\mathscr{H}_{0}$, the Born probability of outcome $z$ given by the pure state differs from the ensemble average by at most $10^{100-0.5\times 10^{20}}\approx 3^{-10^{20}}$, which is close to zero.
+$\Bigl{|}\langle\Psi|\,B\,|\Psi\rangle-\operatorname{tr}(\rho_{0}\,B)\Bigr{|}\leq\frac{\|B\|}{\sqrt{\varepsilon d_{0}}}\,,$ (2)
+where the operator norm $\|B\|$ means the largest absolute eigenvalue of $B$.
+Moreover, for every POVM $\{E_{1},\ldots,E_{L}\}$ and every $\varepsilon>0$, for $(1-\varepsilon)$-most $\Psi\in\mathbb{S}(\mathscr{H}_{0})$, the distribution $\langle\Psi|E_{z}|\Psi\rangle$ is close to the distribution $\operatorname{tr}(\rho_{0}\,E_{z})$ in the sense that for each $z$,
+$\Bigl{|}\langle\Psi|\,E_{z}\,|\Psi\rangle-\operatorname{tr}(\rho_{0}\,E_{z})\Bigr{|}\leq\sqrt{\frac{L}{\varepsilon d_{0}}}\,.$ (3)
+Note that the right-hand side of (3) is small whenever
+$d_{0}\gg\frac{L}{\varepsilon}\,.$ (4)
+Theorem 1 is a particular case of Theorem 3 proven by Teufel et al.
+(2023).
+It can also be derived from a theorem proven by Reimann (2007).
+The statement involving Eq.
+(3) follows from the one involving (2).
+[^6] Here is how: Replace $B$ by $E_{z}$ and $\varepsilon$ by $\varepsilon/L$ in the statement involving (2) and use $\|E_{z}\|\leq 1$.
+Thus, for any $z$, the set of $\Psi$’s satisfying (3) has measure at least $1-\varepsilon/L$, so the intersection of these sets has measure at least $1-\varepsilon$.
+It can be shown that $\operatorname{tr}(\rho_{0}E_{z})$ is exactly the average of $\langle\Psi|E_{z}|\Psi\rangle$ over $\mathbb{S}(\mathscr{H}_{0})$ (using the uniform distribution $u_{0}$); thus, (3) expresses that $\langle\Psi|E_{z}|\Psi\rangle$ is close to its average value whenever (4) holds.
+Theorem 1 is an example of the “concentration-of-measure phenomenon.” Here, we have that for any observation represented by a POVM, typical quantum states in a high-dimensional Hilbert space have very nearly identical Born probability for that observation.
+Let us clarify and elaborate on this theorem.
+Example 1.
+Which size should we consider for $d_{0}$ and $\varepsilon$ in practice?
+For a macroscopic quantum system with $N$ particles, the dimension of an energy shell is usually of order $10^{N}$ (or perhaps more like $10^{1000\,N}$).
+While the dimensions of subspaces corresponding to different macro states differ by huge factors, even the comparatively small subspaces of this kind, corresponding to states far from thermal equilibrium, have dimension of order roughly $10^{N}$.
+For the universe, a widespread estimate is $N=10^{80}$, which motivates us to consider,
+[…]
+Let us turn to $\varepsilon$. Borel (1962) famously argued that an event with probability,
+[…]
+or smaller can be assumed to never occur in the history of the universe.
+Of course, we can make sure that some event of smaller probability occurs, for example by tossing a fair coin 1000 times; then the actual sequence of heads and tails had probability $2^{-1000}\approx 10^{-300}$ like every sequence; but Borel meant that any event specified in advance with probability $\leq 10^{-200}$ is not going to occur.
+Correspondingly, for this value of $\varepsilon$ and any given POVM, we can expect $\Psi_{t_{0}}$ to lie in the set of $\Psi$’s satisfying (3).
+While our results do not require the values (5), (6), we will use them for illustration.
+So for our purposes, (4) is satisfied, for example for $L=2$ or $L=1000$, with enormous margins.
+Our results in §3 will be based on (2), as it provides the best bound for a given outcome $z$.
+Example 2.
+[…]
 Minimal Assumptions.
 Theorem 1 holds in great generality.
-First, it applies to every POVM.
-By the Main Theorem about POVMs, the theorem covers every observational outcome arising from an arbitrary quantum physical experiment.
+First, it applies to an arbitrary POVM.
+(In fact, it even applies to every bounded operator.) By the Main Theorem about POVMs, the theorem covers every possible outcome arising from an arbitrary quantum physical experiment.
 Second, it does not make assumptions about the interaction between the subsystem and the environment.
 Third, it does not invoke chaos, ergodicity, or mixing.
-The typicality in distribution typicality comes from the typicality of quantum states in $\mathbb{S}(\mathscr{H}_{0})$, and the large number comes from $d_{0}$, the large dimension of $\mathscr{H}_{0}$.
+The typicality in distribution typicality comes from the typicality of quantum states in $\mathbb{S}(\mathscr{H}_{0})$ for a given observation, and the large number comes from $d_{0}$, the large dimension of $\mathscr{H}_{0}$.
 Finally, distribution typicality applies to thermal equilibrium and non-equilibrium.
@@
 We do not have distribution typicality in classical mechanics.
-The prediction of any observation from a particular micro-state in the classical phase space has a trivial probability of 00 or $1$, while the expectation values for many observables with respect to a uniform probability distribution (the ensemble averages) are often non-trivial.
-It is false that for every experimental outcome, typical micro-states will assign very nearly the same probability distribution.
+The prediction of any observation from a particular micro-state in the classical phase space has a trivial probability of 00 or $1$, while for many observables their probability distribution induced by the uniform probability distribution in a region of phase space (the ensemble statistics) is non-trivial.
 For example, we can design a coin-flip experiment where half the microstates assign probability 00 and the other half assign probability $1$ to the “Tails” outcome.
 The relevant way in which quantum mechanics differs from classical mechanics is that if a wave function $\Psi=\Psi_{t_{0}}$ gets chosen randomly, then the outcome $Z$ of an observation is “doubly random”: Given $\Psi$, it is random with the Born distribution $\langle\Psi|E_{z}|\Psi\rangle$, and in addition $\Psi$ is itself random.
-The crucial point is that in our situation, different $\Psi$s have very similar superposition weights, with the consequence that they have nearly the same $\langle\Psi|E_{z}|\Psi\rangle$.
+The crucial point is that in our situation, different $\Psi$s have very similar superposition weights, with the consequence that they have nearly identical $\langle\Psi|E_{z}|\Psi\rangle$.
 2.2 Main Motivation: The Past Hypothesis
@@
 115), (Penrose 1979, Lebowitz 1993, Albert 2000, Goldstein et al.
-2020) that the initial state of the universe must be restricted to a set of low-entropy states; in a quantum theory, such a set would be given by a suitable subspace $\mathscr{H}_{0}$ of the Hilbert space $\mathscr{H}$ of the universe, and the presumed additional law might then be formulated as follows:
+2020, Chen 2021) that the initial state of the universe must be restricted to a set of low-entropy states; in a quantum theory, such a set would be given by a suitable subspace $\mathscr{H}_{0}$ of the Hilbert space $\mathscr{H}$ of the universe, and the presumed additional law might then be formulated as follows:
 Past Hypothesis (PH)
@@
 Then the PH requires $\Psi_{t_{0}}$ to have low entropy and to be far from thermal equilibrium.
-Even though $\mathscr{H}_{0}$ has much lower dimension than $\mathscr{H}_{eq}$, $\mathscr{H}_{0}$ is still, like all subspaces representing macro-states, a high-dimensional subspace.
-Realistically, we can expect $d_{0}>>10^{10^{20}}$.
 With respect to the normalized uniform measure $u_{0}$, we expect that typical wave functions in $\mathbb{S}(\mathscr{H}_{0})$ will evolve in a way such that $S_{qB}$ increases in the medium and the long run, satisfying the Developmental Conjecture formulated by Goldstein et al.
 (2020), a version of the Second Law of Thermodynamics for a quantum universe.
-Hence, if we choose (say) $\varepsilon=10^{-200}$, then for any $E_{z}$, typical individual wave functions in $\mathscr{H}_{0}$ will all lead to nearly the same probability distribution as $\rho_{0}$.
 In the philosophical literature, $\rho_{0}$ has been called the Wentaculus density matrix, corresponding to the initial state of the Wentaculus theory (Chen 2021, 2024a).
@@
 3.1 General Distinctions
-There are different kinds of observational indistinguishability (OI) discussed in the philosophical literature.
-We draw some general distinctions regarding (1) OI and empirical equivalence, and (2) in-principle and in-practice OI.
-Empirical equivalence is a relation holding among certain physical theories.
-Two theories may be regarded as empirically equivalent when they make the same predictions about all possible observations.
-For example, one can explore whether BM is empirically equivalent to OQM, GRW, or Everettian quantum mechanics by checking whether BM makes the same empirical predictions as the latter.
-In contrast, OI is primarily a relation holding among models in the same theory.
-In the philosophical literature, there is a large literature about the OI of certain models of general relativity.
-Manchak (2009) showed that for almost every general relativistic spacetime, we can construct another one in such a way that no observation will distinguish between the two spacetimes.
-In this paper, we focus on OI as it is applied to models of quantum theory.
-Here, we shall take an interpretation-neutral approach, and discuss how to apply the result in different interpretations in §3.2.
-Nevertheless, what we show may also be relevant to the issue of empirical equivalence among quantum theories.
-The second distinction is between in-principle OI and in-practice OI.
-People often focus on exact equivalence of predictions (or retrodictions) about outcomes of observations.
-This is done in the general relativistic case in terms of exact agreement (up to an isometry transformation) of past lightcones of observers, and in quantum foundations in terms of exact equality of theoretical distributions.
-In many cases, exact equivalence is sufficient but not necessary for OI.
-We never did and never will probe the exact microscopic detail of the past lightcone.
-In practice we will quickly exhaust our resources long before reaching the in-principle limits.
-Nevertheless, to prove sharp mathematical results, it is sometimes easier to focus on in-principle limits.
-Of course, in-principle OI implies in-practice OI.
-Observation typicality concerns the stronger sense of OI—in principle and not just in practice.
-We show that in-principle OI follows from approximate equality of probabilistic predictions.
-The approximate equality of predictions is to be distinguished from the approximation of frequency to chance.
-For example in Bohmian mechanics, we justify the quantum equilibrium hypothesis using a law-of-large-number argument (Dürr et al.
-1992), with frequency converging to chance: typically (with respect to the quantum equilibrium distribution), the empirical distribution converges to the theoretical distribution.
-The result establishes exact equality of the chance functions (theoretical distributions).
-While exact equality of probabilities is sufficient for in-principle OI, our argument below shows that it is not necessary.
-According to a standard criterion (Nielsen and Chuang 2010, p.86), two quantum states are perfectly distinguishable by some experiment if and only if they are orthogonal.
-If we do not insist on perfect distinguishability, we may say that two quantum states are reliably distinguishable by some experiment if and only if they are (at least) approximately orthogonal.
-We can specialize the criterion for OI to a particular observation $E_{z}$.
-For two quantum states $\Psi_{1}$ and $\Psi_{2}$, we say that they are perfectly distinguishable by observing the outcome $Z=z$ if and only if $\langle\Psi_{1}|\,E_{z}\,|\Psi_{1}\rangle=0$ and $\langle\Psi_{2}|\,E_{z}\,|\Psi_{2}\rangle=1$ or vice versa; let us assume the former is the case.
-If we know that the actual quantum state is either $\Psi_{1}$ or $\Psi_{2}$, the outcome $Z=z$ is a perfect indicator of $\Psi_{2}$.
-The indicator is perfect in the sense that, in 100% of the worlds compatible with the outcome $Z=z$, it correctly indicates the actual quantum state.
-If we do not insist on perfect distinguishability, we can allow the probabilities to be approximately zero or one, so that they are reliably distinguishable by $E_{z}$ just in case $\langle\Psi_{1}|\,E_{z}\,|\Psi_{1}\rangle\approx 0$ and $\langle\Psi_{2}|\,E_{z}\,|\Psi_{2}\rangle\approx 1$, or $\langle\Psi_{1}|\,E_{z}\,|\Psi_{1}\rangle\approx 1$ and $\langle\Psi_{2}|\,E_{z}\,|\Psi_{2}\rangle\approx 0$, i.e., the absolute difference in their probabilities is close to 1.
-The outcome $Z=z$ is a reliable indicator of the quantum state in the sense that, in nearly 100% of the worlds compatible with the outcome $Z=z$, it correctly indicates the actual quantum state.
-This suggests the following sufficient condition:
-OI with respect to $E_{z}$
-If the Born probability distributions of $E_{z}$ assigned by a set of quantum states are all within (say) $10^{-200}$ of each other, then observing the outcome $Z=z$ does not reliably distinguish among those states.
-In what follows, unless specified otherwise, we shall understand OI with this criterion.
-3.2 Observational Indistinguishability of Typical Quantum States
-In this section, we show that distribution typicality implies observation typicality.
-We do this separately for OQM, BM, and GRW.
-The same arguments as for OQM and BM also provide OI for Everettian quantum mechanics.
-(1) OI in OQM.
-Suppose we carry out any experiment $\mathscr{E}$; suppose $\mathscr{E}$ is associated, according to the Main Theorem about POVMs, with the POVM $E$; suppose we obtained the outcome $Z=z$ associated with the operator $E_{z}$.
-Then typical quantum states in a high-dimensional Hilbert space $\mathscr{H}_{0}$ (with $d_{0}>10^{1000}$) are not distinguished by this outcome from $\rho_{0}$.
-Indeed, consider $\mathscr{H}_{0}$ with $d_{0}>10^{1000}$ and $\varepsilon=10^{-200}$.
-By Theorem 1 for any POVM element $E_{z}$, for ($1-\varepsilon)$-most $\Psi$ in $\mathscr{H}_{0}$, $|\langle\Psi|\,E_{z}\,|\Psi\rangle-\operatorname{tr}(\rho_{0}E_{z})|\leq\frac{1}{\sqrt{\varepsilon d_{0}}}<10^{-400}$.
-Hence, the Born probability distributions of $E_{z}$ assigned by ($1-\varepsilon)$-most $\Psi$ are all within $\tfrac{1}{2}\varepsilon$ of that of $\rho_{0}$, and thus within $\varepsilon$ of each other.
-As discussed above, a reliable distinction would require the observation of an event that has probability near 1 in one case and near 0 in the other, which does not happen here.
-Therefore, observing $E_{z}$ does not reliably distinguish those states from each other or from the normalized projection.
-So, we arrive at observation typicality in $\mathscr{H}_{0}$.
-Again, we may think of $\mathscr{H}_{0}$ as the subspace demanded by the Past Hypothesis.
-Thus, for any observation $E_{z}$, typical quantum states in $\mathscr{H}_{0}$ are not distinguished by $E_{z}$ from the Wentaculus density matrix.
-Neither does it distinguish typical quantum states in $\mathscr{H}_{0}$ from each other.
-In quantum theory with unitary evolution, this extends to later states too: Since typical $\Psi_{t}$ from the appropriately evolved subspace $U_{t}\mathscr{H}_{0}$ lead to nearly the same probability for outcome $Z=z$, the observation of this outcome does not distinguish among different $\Psi_{t}$.
-(2) OI in BM. In BM, a key fact about OI is this:
-the probability distribution of the configuration $Q_{t}$ of the universe at time $t$, i.e., the $|\Psi_{t}|^{2}$ distribution in configuration space, is nearly independent of $\Psi_{t_{0}}$ for typical $\Psi_{t_{0}}\in\mathbb{S}(\mathscr{H}_{0})$.
-(5)
-This follows from Theorem 1 by taking the POVM $E$ to be the Heisenberg-evolved configuration observable, i.e., the projection-valued measure jointly diagonalizing all position operators at time $t$, or
-$E(B)=U_{t}^{\dagger}1_{B}U_{t},$ (6)
-where $B$ is any subset of configuration space and $1_{B}$ the multiplication operator multiplying by the characteristic function of $B$.
-Note that the probability distribution associated with this POVM is exactly the $|\Psi_{t}|^{2}$ distribution:
+There are various kinds of observational indistinguishability (OI).
+Here we clarify the distinction between in-principle OI and in-practice OI.
+Discussions in physics often focus on exact equivalence of predictions (or retrodictions) about observational outcomes.
+In general relativity, this is framed in terms of exact agreement (up to an isometry transformation) of the past lightcones of observers (Manchak 2009).
+In quantum foundations, many results pertain to exact equality of theoretical probability distributions.
+While exact equivalence is often sufficient for OI, it is not always necessary.
+Resource limitations mean that we sometimes reach in-practice OI long before hitting the in-principle limits.
+Nevertheless, sharp mathematical results are sometimes more easily proven by focusing on in-principle limits.
+Naturally, in-principle OI implies in-practice OI.
+Observation typicality pertains to the stronger sense of OI—in principle, rather than merely in practice: No progress in technology would improve the situation, nor would more costly experiments or any other added resources.
+That is because we show that no substantial information about $\Psi_{t_{0}}$ can be obtained from any experiment using any resources, whenever the dimension $d_{0}$ of $\mathscr{H}_{0}$ is sufficiently large.
+It may seem surprising that an in-principle limitation can follow from an approximate equality (such as $\langle\Psi|E_{z}|\Psi\rangle\approx\operatorname{tr}(\rho_{0}E_{z})$), for two reasons.
+First, it would seem that nearby quantities can be distinguished better by measuring them with higher resolution; but here, the quantities are, not measurement outcomes, but probabilities of outcomes.
+Second, it would seem that probabilities can be measured more accurately if we repeat the experiment sufficiently often; but here, that is not possible, as we will discuss in §3.5.
+3.2 Mathematical Descriptions
+In §3.3, we will discuss for OQM, BM, GRW, and EQM which limitations for observers arise and how. But first, in this section, we describe the mathematical facts on which theses conclusions rest.
+3.2.1 Not Reliably Distinguishable
+We first show that, as a consequence of Theorem 1, typical vectors in $\mathscr{H}_{0}$ are not reliably distinguishable.
+Let us make precise what we mean by “reliable.” Suppose we use outcome $z$ to determine whether $A$ is the case or $B$ is the case.
+We say that $E_{z}$$(1-\delta)$-reliably” distinguishes $A$ from $B$ just in case $|\mathbb{P}_{A}(Z=z)-\mathbb{P}_{B}(Z=z)|\geq 1-\delta$, where $\mathbb{P}_{A}$ means the probability if $A$ is the case, and likewise for $\mathbb{P}_{B}$.
+Thus, a reliability of 95% requires that the two probabilities differ by at least $95\%$.
+The reliability lies between 0 and 1, where 1 means perfectly reliable and 0 means completely unreliable.
+[^8] Under the decision rule “if $z$ occurs, guess $A$; otherwise guess $B$,” and assume $\mathbb{P}_{A}(z)\geq\mathbb{P}_{B}(z)$.
+Then $|\mathbb{P}_{A}(z)-\mathbb{P}_{B}(z)|\geq 1-\delta$ implies $\text{Sensitivity}=\mathbb{P}_{A}(z)\geq 1-\delta,$ and $\text{Specificity}=1-\mathbb{P}_{B}(z)\geq 1-\delta,$ so the Type I error (false-positive) rate $\mathbb{P}_{B}(z)\leq\delta$ and the Type II error (false-negative) rate $1-\mathbb{P}_{A}(z)\leq\delta$.
+Hence, smaller $\delta$ means smaller Type I and Type II errors in hypothesis testing.
+Definition 1.
+Let $\{E_{z}:z\in\mathscr{Z}\}$ be any POVM, $z$ any element of $\mathscr{Z}$, and $\delta>0$.
+We say $E_{z}$ $(1-\delta)$-reliably distinguishes the density matrices $\rho_{A},\rho_{B}$ if and only if $|\operatorname{tr}(\rho_{A}E_{z})-\operatorname{tr}(\rho_{B}E_{z})|\geq 1-\delta$.
+Note that distinguishing at a reliability level $1-\delta$ also counts as distinguishing at any lower reliability level (i.e., larger $\delta$).
+Example 3.
+Consider again Example 2, with $E_{z}=P_{\uparrow}=\frac{1}{2}(I+\sigma_{z})$.
+If $\delta$ is taken to be $0.1,0.2,0.3$, then $E_{z}$ $(1-\delta)$-reliably distinguishes, respectively, $1\%,4\%,9\%$ of the total pairs of $\psi$’s in $\mathbb{C}^{2}$.
+(These fractions are given by $\delta^{2}$.
+[^9] Proof: Let $p_{A}=\langle\Psi_{A}|P_{\uparrow}|\Psi_{A}\rangle$ and likewise $p_{B}$.
+The set of $(p_{A},p_{B})\in[0,1]^{2}$ with $|p_{A}-p_{B}|\geq 1-\delta$ is the union of the triangles $(0,1-\delta)-(\delta,1)-(0,1)$ and $(1-\delta,0)-(1,\delta)-(1,0)$, which has area $\delta^{2}$.
+By Footnote 7, the set of $(\Psi_{A},\Psi_{B})$ with such $(p_{A},p_{B}$) has the same measure.)
+We can show that for fixed $\delta$, in higher dimensional Hilbert spaces, Theorem 1 forces the fraction of pairs that $E_{z}$ $(1-\delta)$-reliably distinguishes to drop quickly and become essentially zero.
+Corollary 1.
+Let $\{E_{z}:z\in\mathscr{Z}\}$ be any POVM, $z$ any element of $\mathscr{Z}$, and $\varepsilon>0$.
+If $d_{0}>4/\varepsilon$, then for $(1-\varepsilon)$-most $\Psi_{A}\in\mathbb{S}(\mathscr{H}_{0})$, $E_{z}$ fails to $\frac{1}{2}$-reliably distinguish $\rho_{A}:=|\Psi_{A}\rangle\langle\Psi_{A}|$ from $\rho_{B}:=\rho_{0}$.
+Furthermore, If $d_{0}>16/\varepsilon$, then for $(1-\varepsilon)^{2}$-most choices of $\Psi_{A},\Psi_{B}$, $E_{z}$ fails to $\frac{1}{2}$-reliably distinguish $\rho_{A}$ from $\rho_{B}:=|\Psi_{B}\rangle\langle\Psi_{B}|$.
+This follows from Theorem 1 via a few steps of algebra.
+[^10] Proof of the first statement: By (2), $|\operatorname{tr}(\rho_{A}E_{z})-\operatorname{tr}(\rho_{B}E_{z})|=|\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle-\operatorname{tr}(\rho_{0}E_{z})|\leq\|E_{z}\|/\sqrt{\varepsilon d_{0}}=1/\sqrt{\varepsilon d_{0}}<1/2$.
+Second statement: $\Psi_{A}$ and $\Psi_{B}$ are chosen like independent random variables, so for $(1-\varepsilon)^{2}$-most pairs $(\Psi_{A},\Psi_{B})$, (2) for $B=E_{z}$ holds for both $\Psi$s, and thus $|\operatorname{tr}(\rho_{A}E_{z})-\operatorname{tr}(\rho_{B}E_{z})|=|\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle-\langle\Psi_{B}|E_{z}|\Psi_{B}\rangle|\leq|\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle-\operatorname{tr}(\rho_{0}E_{z})|+|\langle\Psi_{B}|E_{z}|\Psi_{B}\rangle-\operatorname{tr}(\rho_{0}E_{z})|\leq 2/\sqrt{\varepsilon d_{0}}<1/2$.
+Note that it trivially follows that $\rho_{A},\rho_{B}$ cannot be $(1-\delta)$-reliably distinguished for any $\delta$ between 0 and 1/2.
+For example, for dimension $d_{0}=10^{100}$ and any $20<\delta<1/$ 2, $E_{z}$ can $(1-\delta)$-reliably distinguish at most the fraction $3.2\times 10^{-99}$ of the pairs of wave functions.
+Corollary 1 should be contrasted with the statement that for large $d_{0}$ and purely randomly and independently chosen $\Psi_{A},\Psi_{B}\in\mathbb{S}(\mathscr{H}_{0})$, the two vectors will be nearly orthogonal, so there exists a POVM that distinguishes them reliably.
+Note, however, that if we keep this POVM $E$ fixed and consider new, random $\Psi_{A},\Psi_{B}$, then the two yield nearly the same Born distribution for $E$.
+So, the two statements differ in the order of quantifiers (“for most,” “there is”): for most $\Psi_{A},\Psi_{B}$, there is $E$ that reliably distinguishes $\Psi_{A},\Psi_{B}$; but according to Corollary 1, it is not the case that there is $E$ such that for most $\Psi_{A},\Psi_{B}$, $E$ reliably distinguishes $\Psi_{A},\Psi_{B}$.
+Likewise, for every given $\Psi_{A}\in\mathbb{S}(\mathscr{H}_{0})$, there is a POVM $E$ that distinguishes $|\Psi_{A}\rangle\langle\Psi_{A}|$ $(1-d_{0}^{-1})$-reliably from $\rho_{0}$: just take $E_{A}=|\Psi_{A}\rangle\langle\Psi_{A}|$ and $E_{B}=I-E_{A}$.
+[^11] Proof: $\operatorname{tr}(\rho_{A}E_{A})=1$ and $\operatorname{tr}(\rho_{0}E_{B})=1-\operatorname{tr}(\rho_{0}E_{A})=1-\langle\Psi_{A}|\rho_{0}|\Psi_{A}\rangle=1-d_{0}^{-1}$.
+Corollary 1 should also be contrasted with the problem of distinguishing between the case of a mixed state $\rho_{0}$ and the case of a random pure state $\Psi$ with distribution $u_{0}$; see §3.5.2 for details.
+3.2.2 Not Even Unreliably Distinguishable
+Next, we show that typical vectors in $\mathscr{H}_{0}$ are not even unreliably distinguishable if the observation isn’t too unlikely.
+Consider an experiment with POVM $\{E_{z}:z\in\mathscr{Z}\}$ and outcome $Z\in\mathscr{Z}$, and consider two density matrices $\rho_{A},\rho_{B}$. Suppose that for a particular $z$,
+$\mathbb{P}_{A}(Z=z)=\operatorname{tr}(\rho_{A}E_{z})=\frac{3}{4}~{}~{}~{}\text{and}~{}~{}~{}\mathbb{P}_{B}(Z=z)=\operatorname{tr}(\rho_{B}E_{z})=\frac{1}{4}.$ (8)
+Suppose we observed the outcome $z$; then, although it does not distinguish reliably between $\rho_{A}$ and $\rho_{B}$ at a reliability level close to 1, one can argue that this outcome at least gives us some probabilistic information, as it favors option $A$ over option $B$ by a factor of 3.
+We now investigate how strongly our observations can favor $\Psi_{A}$.
+Definition 2.
+For any density matrices $\rho_{A},\rho_{B}$ on $\mathscr{H}$ and any POVM $\{E_{z}:z\in\mathscr{Z}\}$, we define the favoring ratio for $A$ given $z$ to be
+$r_{A|z}:=\frac{\operatorname{tr}(\rho_{A}E_{z})}{\operatorname{tr}(\rho_{B}E_{z})}\,.$ (9)
+The next statement expresses that in our case, the favoring ratio typically deviates from 1 by merely a negligible amount.
+Corollary 2.
+Let $\{E_{z}:z\in\mathscr{Z}\}$ be any POVM, $z$ any element of $\mathscr{Z}$, and $\varepsilon>0$.
+If $\operatorname{tr}(\rho_{0}E_{z})>\varepsilon$, then for $(1-\varepsilon)$-most $\Psi_{A}\in\mathbb{S}(\mathscr{H}_{0})$, the favoring ratio of $\rho_{A}=|\Psi_{A}\rangle\langle\Psi_{A}|$ over $\rho_{B}=\rho_{0}$ obeys
+$1-\frac{1}{\sqrt{\varepsilon^{3}d_{0}}}\leq r_{A|z}\leq 1+\frac{1}{\sqrt{\varepsilon^{3}d_{0}}}\,.$ (10)
+Furthermore, if also $d_{0}>9\varepsilon^{-3}$, then for $(1-\varepsilon)^{2}$-most choices of $\Psi_{A},\Psi_{B}$ in $\mathbb{S}(\mathscr{H}_{0})$, the favoring ratio of $\rho_{A}=|\Psi_{A}\rangle\langle\Psi_{A}|$ over $\rho_{B}=|\Psi_{B}\rangle\langle\Psi_{B}|$ obeys
+$1-\frac{2}{\sqrt{\varepsilon^{3}d_{0}}}\leq r_{A|z}\leq 1+\frac{3}{\sqrt{\varepsilon^{3}d_{0}}}\,.$ (11)
+In both cases, the favoring ratio is close to 1 whenever $d_{0}\gg\varepsilon^{-3}$.
+This follows from Theorem 1 through some steps of algebra.
+[^12] Proof: By hypothesis, $\eta:=(\varepsilon d_{0})^{-1/2}/\operatorname{tr}(\rho_{0}E_{z})\leq 1/\sqrt{\varepsilon^{3}d_{0}}$.
+By Theorem 1, $(1-\varepsilon)$-most $\Psi_{A}$ obey (2) with $B=E_{z}$; dividing by $\operatorname{tr}(\rho_{0}E_{z})$, we obtain that $\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle/\operatorname{tr}(\rho_{0}E_{z})\in[1-\eta,1+\eta]$.
+Second statement: By hypothesis, $\eta<1/3$.
+By Theorem 1, for $(1-\varepsilon)^{2}$-most pairs, both $\Psi$s obey (2) with $B=E_{z}$; we obtain the same relation as stated above for $\Psi_{A}$ also for $\Psi_{B}$.
+If $\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle\geq\langle\Psi_{B}|E_{z}|\Psi_{B}\rangle$, then $1\leq\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle/\langle\Psi_{B}|E_{z}|\Psi_{B}\rangle\leq(1+\eta)/(1-\eta)=1+2\eta/(1-\eta)<1+3\eta$ using $\eta<1/3$.
+If $\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle\leq\langle\Psi_{B}|E_{z}|\Psi_{B}\rangle$, then $1\geq\langle\Psi_{A}|E_{z}|\Psi_{A}\rangle/\langle\Psi_{B}|E_{z}|\Psi_{B}\rangle\geq(1-\eta)/(1+\eta)=1-2\eta/(1+\eta)\geq 1-2\eta$.
+In words, it means that the outcome $z$, if it is not very improbable, leads to no significant probabilistic preference of $\Psi_{A}$ over $\rho_{0}$, or of $\rho_{0}$ over $\Psi_{A}$, or of $\Psi_{A}$ over $\Psi_{B}$, or of $\Psi_{B}$ over $\Psi_{A}$.
+What about outcomes that are very improbable?
+They should not occur if (as we assume) the theory is correct, unless there are so many possible outcomes that the probabilities of the very improbable ones can add up to a significant value.
+That situation occurs in the example of 1000 coin tosses mentioned already near (6), so, unlike the situation considered by Borel, we need to also consider here events not specified in advance.
+That is, we need to reconsider the relevant size of $\varepsilon$ because $\varepsilon$ now appears in two roles: the maximal measure of the exceptional $\Psi$’s and the minimal probability of the improbable observations.
+So, how small could the probability of an observed outcome be?
+Like the coin toss example, an experiment can have a large number $L$ of possible outcomes, and if all have equal or similar probability, then these probabilities will be tiny of order $1/L$.
+If we need to distinguish between $L$ different observations, we can perhaps reach $L=10^{10^{10}}$ by collecting $10^{9}$ data points, each given by a number with 10 decimal digits.
+On the other hand, as of this writing (2025), the total storage capacity of all computers in the world is estimated at about $10^{26}$ bytes, so we would not be able to even record the outcome if $L$ exceeded $10^{10^{30}}$, which motivates us to consider,
 […]
-for any subset $B$ of configuration space.
-Now $Q_{t}$ comprises the exact positions of all particles in the universe, so observers inside a Bohmian universe cannot possibly measure $Q_{t}$; at best, they can measure a very coarse-grained version of $Q_{t}$.
-But the point is that even complete information about $Q_{t}$ would not help with distinguishing between typical $\Psi_{t_{0}}$ and $\rho_{0}$ (or between two typical wave functions); and a fortiori, a coarse-grained version does not help.
-The essence of the difficulty with distinguishing is an instance of the general problem of deciding between two probability distributions $f_{1}$ and $f_{2}$ (say, on $\mathbb{R}^{n}$) after we are given just one point $X$ that was randomly chosen with distribution either $f_{1}$ or $f_{2}$.
-If $f_{1}$ and $f_{2}$ are disjoint, then this is possible; if they are nearly disjoint, then it is still possible quite reliably (for example, at the confidence level of $95\%$ often used in statistics).
-But if $f_{1}$ and $f_{2}$ overlap significantly, and if the observed $X$ lies in the overlap region, then this value of $X$ could have arisen from either one, so we cannot decide which $f$ was used.
-A reliable decision is possible only if $f_{1}(X)\approx 0$ and $f_{2}(X)$ is significantly non-zero or vice versa.
-Now in our case, $f_{i}=|\Psi_{i}|^{2}$ and $f_{1}\approx f_{2}$ by (5), so they overlap almost exactly and everywhere; values of $X=Q_{t}$ with $f_{1}(X)\approx 0$ will also have $f_{2}(X)\approx 0$ and therefore, with overwhelming probability, are not going to occur for either $f_{1}$ or $f_{2}$.
-So, the two cannot be distinguished at any time $t$ (not reliably and in fact, as we will discuss in §3.4, not even unreliably).
-This conclusion covers also the possibility that observers might make observations (say, with telescopes) or experiments, as the outcomes of these observations or experiments will be recorded in the configuration of some particles (e.g., when scientists publish their findings).
-It also covers the possibility of making observations or experiments at several times $t_{1},t_{2},\ldots,t_{k}$, since the outcomes will be recorded and thus can be read off from $Q_{t}$ at $t\geq t_{k}$.
-On the other hand, the precise microscopic trajectories of all Bohmian particles may contain more information about $\Psi_{t_{0}}$, but it is known (Dürr et al.
-1992) that observers inside a Bohmian universe cannot measure positions at several times without introducing decoherence and thereby changing the trajectory.
-Therefore, they do not have access to the full trajectories, and the observational indistinguishability just derived remains valid.
-In other words, there is a known limitation to knowledge about full trajectories—the absolute uncertainty in Bohmian mechanics.
-(3) OI in GRWf. We consider GRWf, the GRW theory with a flash ontology in spacetime (e.g., Tumulka 2022b, sec. 3.3.4). A key fact about OI in GRWf is this:
-the joint probability distribution of all flashes in the universe up to time $t$ is nearly independent of $\Psi_{t_{0}}$ for typical $\Psi_{t_{0}}\in\mathbb{S}(\mathscr{H}_{0})$. (8)
-This follows from Theorem 1 by taking the POVM $E$ to be the POVM governing the joint distribution of all flashes (Tumulka 2022b, sec.
-5.1.1 and 7.8).
-As in BM, observers in a GRWf universe do not have access to the full pattern of flashes, only to a coarse-grained version of it.
-But even the exact pattern would not provide the information needed for distinguishing a typical $\Psi_{t_{0}}$ from $\rho_{0}$ or from another typical initial wave function, for the same reasons based on Theorem 1 as discussed for BM above.
-A difference to the Bohmian case is that the exact history of the universal configuration in BM may provide sufficient information to determine $\Psi_{t_{0}}$, whereas in GRWf even the exact history of all flashes is of no help.
-Thus, GRWf provides, in a sense, a stronger kind of limitation to knowledge than BM.
-Another difference to BM is that inhabitants of a GRWf universe living at time $t$ may very well find out a lot about $\Psi_{t}$.
-After all, $\Psi_{t}$ is not unitarily evolved from $\Psi_{t_{0}}$ but has collapsed; for example, if we find Schrödinger’s cat alive then its wave function will have collapsed to (approximately) the wave function of a live cat.
-(Note that this still does not give us new information about $\Psi_{t_{0}}$.)
-3.3 Remarks
-Let us return to the general result of observation typicality and discuss the order of quantifiers.
-It has the form “for any $E$ and for most $\Psi$, $s(\Psi,E)$.” This is weaker than the statement “for most $\Psi$ and for any $E$, $s(\Psi,E)$,” which is false in this context.
-It is easy to confuse the two.
-The false but stronger statement corresponds to the following:
-Super-Strong Observation Typicality
-For most quantum states in a high-dimensional Hilbert subspace $\mathscr{H}_{0}$, for any observation $E_{z}$, they cannot be distinguished by $E_{z}$.
-Super-strong observation typicality implies observation typicality; it is false because if $E_{z}=|\psi\rangle\langle\psi|$ is a 1d projection in the direction of a particular $\psi\in\mathbb{S}(\mathscr{H}_{0})$, then $\mathbb{P}_{\Psi=\psi}(Z=z)=1$, whereas $\mathbb{P}_{\rho_{0}}(Z=z)=1/d_{0}\approx 0$.
-The super-strong property requires a uniformity in the class of typical quantum states, while observation typicality is compatible with the possibility that the typicality class depends on the choice of $E_{z}$.
-However, the weaker statement—observation typicality—is sufficient for our purposes.
-OI of typical quantum states is true for any chosen experiment, so it is true for the actual experiment that we happen to perform at an arbitrary time.
-Another remark concerns alternative arguments for OI using other typicality theorems, specifically canonical typicality (Gemmer et al.
-2009, Popescu et al.
-2006, Goldstein et al.
-2006a) and dynamical typicality as stated in (Teufel et al.
-2023).
-We will discuss their connection to distribution typicality and observation typicality elsewhere (Chen and Tumulka 2024).
-Here we make two brief comments:
-  1. 1.
-    Suppose we have access to only local observables pertaining to a small subsystem $\mathscr{S}$ of the universe (less than half the number of degrees of freedom in the universe).
-Then our observations cannot distinguish typical quantum states $\Psi_{t_{0}}$ of the universe from each other.
-After all, canonical typicality, applied to a subspace $\mathscr{H}_{R}$ now not given by the micro-canonical energy shell (as in most discussions of canonical typicality) but by $\mathscr{H}_{R}=U_{t}\mathscr{H}_{0}$, asserts that most $\Psi_{t}\in\mathscr{H}_{R}$ have nearly the same reduced density matrix $\rho_{\mathscr{S}}=\operatorname{tr}_{\mathscr{S}^{c}}|\Psi_{t}\rangle\langle\Psi_{t}|$ in $\mathscr{S}$ (where $\mathscr{S}^{c}$ denotes the complement of $\mathscr{S}$).
-But the Born distribution of any observable in $\mathscr{S}$ depends on $\Psi_{t}$ only through $\rho_{\mathscr{S}}$; thus, the Born distribution accessible to observers like us is nearly independent of $\Psi_{t_{0}}$.
-    In this context, the super-strong form of OI is actually true: for a typical quantum state $\Psi_{t_{0}}$, there is no local observation at all that can distinguish it from $\rho_{0}$.
-After all, once $\operatorname{tr}_{\mathscr{S}^{c}}|\Psi_{t}\rangle\langle\Psi_{t}|\approx\operatorname{tr}_{\mathscr{S}^{c}}(U_{t}\rho_{0}U_{t}^{\dagger})$, all observables in $\mathscr{S}$ have nearly the same Born distribution in $\Psi_{t}$ as in $U_{t}\rho_{0}U_{t}^{\dagger}$.
-  2. 2.
-    Dynamical typicality means here the statement that most $\Psi_{t_{0}}\in\mathbb{S}(\mathscr{H}_{0})$ have for every $t$ nearly the same Born distribution for all macroscopic observables (and nearly the same as $\rho_{0}$).
-If any scientist could distinguish $\Psi_{t_{0}}$ from $\rho_{0}$, then she or he could publish the result, say at time $t$, so there would be a macroscopic difference between $\Psi_{t_{0}}$ and $\rho_{0}$, which is impossible by dynamical typicality.
-3.4 A Bayesian Analysis
-So far, we have argued that no observations can reliably distinguish typical universal quantum states from each other.
-What about unreliable tests that may nonetheless yield probabilistic information about the quantum state?
-We can use the previous results to give a Bayesian analysis.
-Here we provide two quantitative bounds of the probabilistic information obtainable from each observation.
-It turns out that insofar as the observation is not too improbable, it will not give us substantial probabilistic information about the universal quantum state.
-Hence, there is a sense that even unreliably, universal quantum states cannot be distinguished.
+Then, since the number of outcomes would realistically be less than $10^{10^{30}}$, the probabilities less than $\varepsilon$ would add up to less than $10^{-9\cdot 10^{30}}$.
+Thus, outcomes with probability less than $\varepsilon$ will not be observed given the computational limitations, and still our examples satisfy $d_{0}=10^{10^{80}}\gg 10^{3\cdot 10^{31}}=\varepsilon^{-3}$.
+3.2.3 A Bayesian Analysis
+Theorem 1 also allows us to perform a Bayesian analysis, i.e., to mathematically express the probabilistic information provided by the observation of the outcome $z$.
+Here, we provide two quantitative bounds on this information.
+They have the consequence that, if the observation is not too improbable, it cannot yield substantial probabilistic information about the universal quantum state.
+This result invites further analysis of its implications for Bayesian epistemology and convergence to the truth (see Lin (2024) for a survey), which we leave for future work.
 Suppose the prior probability distribution over quantum states in $\mathscr{H}_{0}$, $\mathbb{P}(\Psi_{t_{0}})$, is given by the Past Hypothesis, i.e., by $u_{0}$.
@@
 Thus, using (13) again, $\mathbb{P}(\Psi\in S|Z=z)=\int_{S}u_{0}(d\psi)\,f(\psi)=\int_{S\cap M}u_{0}(d\psi)\,f(\psi)+\int_{S\cap M^{c}}u_{0}(d\psi)\,f(\psi)\leq\int_{S\cap M}u_{0}(d\psi)\,(1+\varepsilon)+\int_{M^{c}}u_{0}(d\psi)\,f(\psi)<(1+\varepsilon)\,u_{0}(S\cap M)+2\varepsilon\leq(1+\varepsilon)u_{0}(S)+2\varepsilon\leq u_{0}(S)+3\varepsilon$.
-The assumption $d_{0}>\varepsilon^{-5}$ is reasonable; even when $\varepsilon=10^{-200}$, the condition is still satisfied as $\varepsilon^{-5}=10^{1000}\leq 10^{10^{20}}\leq d_{0}$.
-In this case, the posterior probability distribution differs minimally from the prior and is almost independent of any observation.
-This conclusion requires the assumption that the observation is not too improbable.
-Hence, distribution typicality entails not only OI, but also that, in general, observations reveal very little about the initial quantum state.
-Notice how this fact fails in classical mechanics.
-Observations, even when they are probable, can reveal a great deal about the classical initial state.
-For example, observing the Tails outcome of a classical coin-flip experiment can rule out half the initial conditions, namely those that predict a Heads outcome.
-Observations cannot reveal so much in quantum mechanics because of distribution typicality and observation typicality, two facts whose analogues fail in classical statistical mechanics.
+The assumption $d_{0}>\varepsilon^{-5}$ is mild; for example for (12), $\varepsilon^{-5}=10^{5\cdot 10^{31}}$ is far less than (5).
+The physical consequence of these corollaries will be that, in general, observations reveal very little about the initial quantum state.
+Example 4.
+To illustrate that the assumption of large $d_{0}$ cannot be dropped, we compute $f(\psi)$ explicitly in a case with $d_{0}=2$: Let $\{|b_{1}\rangle,|b_{2}\rangle\}$ be an orthonormal basis of $\mathscr{H}_{0}$, and let $E_{z}=|b_{1}\rangle\langle b_{1}|$.
+Then [^15] Proof: $\mathbb{P}(Z=z|\Psi=\psi)=\bigl{|}\langle\psi|b_{1}\rangle\bigr{|}^{2}$ by Born’s rule, so $\mathbb{P}(Z=z,\Psi\in d\psi)=\bigl{|}\langle\psi|b_{1}\rangle\bigr{|}^{2}u_{0}(d\psi)$ and $\mathbb{P}(Z=z)=1/2$ by symmetry, so $\mathbb{P}(\Psi\in d\psi|Z=z)=2\bigl{|}\langle\psi|b_{1}\rangle\bigr{|}^{2}u_{0}(d\psi)=f(\psi)\,u_{0}(d\psi)$, which yields (15).
+$f(\psi)=2\bigl{|}\langle\psi|b_{1}\rangle\bigr{|}^{2}\,,$ (15)
+which varies between 0 and 2 and deviates significantly from 1 for a substantial fraction of $\psi$’s: indeed, it follows from Example 2 that $f(\psi)$ exceeds 1.8 for 10% of all $\psi$’s.
+Example 5.
+To illustrate that the assumption $\operatorname{tr}(\rho_{0}E_{z})>\varepsilon$ cannot be dropped, we consider an orthonormal basis $\{|b_{1}\rangle,\ldots,|b_{d_{0}}\rangle\}$ of $\mathscr{H}_{0}$ and the POVM $E_{z}=|b_{z}\rangle\langle b_{z}|$ for $z=1,\ldots,d_{0}$.
+Then $\operatorname{tr}(\rho_{0}E_{z})=1/d_{0}$ for any $z$, which is much less than $\varepsilon$ if $d_{0}>\varepsilon^{-5}$.
+By the same reasoning as for (15), $f(\psi)=d_{0}|\langle\psi|b_{z}\rangle|^{2}$, which is far from being constantly 1.161616Proof: It is known that for random $\Psi$ with uniform distribution and large $d_{0}$, $\langle\Psi|b_{z}\rangle$ has an approximately Gaussian distribution in the complex plane with mean 0 and variance $\mathbb{E}[|\langle\Psi|b_{z}\rangle|^{2}]=1/d_{0}$.
+Since this distribution gives measure 1/4 to the disk around the origin with radius $\sqrt{\ln(4/3)/d_{0}}$ and 1/4 to the complement of the disk around the origin with radius $\sqrt{\ln(4)/d_{0}}$, we have that $f(\psi)<\ln(4/3)<0.288$ for 1/4 of all $\psi$’s and $f(\psi)>\ln(4)>1.38$ for 1/4 of all $\psi$’s.
+3.3 Physical Consequences: First Argument for OI
+We now explain how these mathematical facts imply OI in the sense of observation typicality in OQM, BM, GRW theory, and EQM.
+We have two arguments, the first of which will be presented now, the second in §3.4; a comparison of the two will be made in §3.5.1.
+The first argument is based on the Main Theorem about POVMs (see §2.1), which is known to be valid in OQM, BM, GRW theory, and EQM.
+The argument puts together the steps we have laid out already.
+By the Past Hypothesis, $\Psi_{t_{0}}$ is typical in $\mathscr{H}_{0}$, which has dimension $d_{0}\geq 10^{10^{80}}\gg\varepsilon^{-5}$ even for $\varepsilon=10^{-10^{31}}$.
+Suppose we carry out any experiment $\mathscr{E}$; $\mathscr{E}$ is associated, according to the Main Theorem about POVMs, with some POVM $E$; suppose we obtained the outcome $Z=z$ associated with the operator $E_{z}$.
+Then according to Corollary 1, typical quantum states in $\mathscr{H}_{0}$ are not reliably distinguished by this outcome from $\rho_{0}$.
+Suppose further $\operatorname{tr}(\rho_{0}E_{z})>\varepsilon$.
+By Corollary 2, the outcome does not significantly favor one $\Psi$ over another or over $\rho_{0}$.
+And by Corollaries 3 and 4, the Bayesian posterior distribution of $\Psi$ does not significantly differ from the prior.
+Thus, regardless of which experiment observers are making, the outcome $z$ provides no significant information about $\Psi$.
+This completes the argument.
+This reasoning uses only one outcome $z$ and the associated operator $E_{z}$; it does not matter how many other $z^{\prime}$ there are and what the $E_{z^{\prime}}$ are.
+OI of typical quantum states is true for any experiment, so it is true for the actual experiment that we happen to perform at an arbitrary time with $z$ standing in for the actual outcome.
+In quantum theory with unitary evolution, OI extends to later states too: Since typical $\Psi_{t}$ from the appropriately evolved subspace $U_{t}\mathscr{H}_{0}$ lead to nearly identical probability for outcome $Z=z$, the observation of this outcome does not distinguish among different $\Psi_{t}$.
+3.4 Physical Consequences: Second Argument for OI
+Our second argument also works in all of OQM, BM, GRW theory and EQM.
+It is more direct than the first argument, particularly in BM and the flash version of GRW.
+We will formulate it first in BM, and afterwards comment on its translation to the other versions of QM.
+The argument concerns the configuration $Q_{t}$ of the universe at an arbitrary time $t$; $Q_{t}$ is typical with respect to the $|\Psi_{t}|^{2}$ distribution, and this distribution is given in terms of the initial wave function $\Psi_{t_{0}}$ of the universe by a POVM (in fact, a PVM (projection-valued measure)) $P$,
+[…]
+for any subset $\Omega$ of the configuration space of the universe.
+This POVM $P$ is the Heisenberg-evolved configuration observable, i.e., the projection-valued measure jointly diagonalizing all position operators at time $t$, or $P(\Omega)=U_{t}^{\dagger}1_{\Omega}U_{t}$, where $1_{\Omega}$ is the multiplication operator multiplying by the characteristic function of $\Omega$.
+Now $Q_{t}$ comprises the exact positions of all particles in the universe, so observers inside a Bohmian universe cannot possibly measure $Q_{t}$; at best, they can measure a very coarse-grained version of some components of $Q_{t}$.
+But the crucial point here is that any information available to the inhabitants of a Bohmian universe at time $t$ will be recorded in $Q_{t}$, in fact in whether $Q_{t}$ lies in some (suitably coarse, macroscopic) set $\Omega$.
+For example, observations (and conclusions drawn from them) will be reported in books and journals, and thus in the configuration of ink particles.
+(For more detailed discussions, see (Albert 2015, pp.
+83-85).) More precisely, we make use of the following
+Necessary Condition
+If an event is known to inhabitants of a Bohmian universe at time $t$ then it is of the form $Q_{t}\in\Omega$ for some set $\Omega$.
+The argument for OI in BM consists of applying Corollaries 14 to the POVM $\{E_{A},E_{B}\}=\{P(\Omega),P(\Omega^{c})\}$ for some subset $\Omega$ of the configuration space $\mathbb{R}^{3N}$ of the universe and its complement $\Omega^{c}=\mathbb{R}^{3N}\setminus\Omega$.
+As a consequence of the necessary condition just mentioned and Corollaries 14, no event known to the inhabitants can reliably distinguish $\Psi_{A}$ from $\rho_{0}$, or can significantly favor $\Psi_{A}$ over $\rho_{0}$ or $\Psi_{B}$, or can substantially alter the Bayesian posterior distribution of $\Psi_{t_{0}}$ (for the last two disjuncts provided $\operatorname{tr}(\rho_{0}P(\Omega))>\varepsilon$, i.e., the event $Q_{t}\in\Omega$ is not too improbable on average).
+Thus, at no time $t$ can inhabitants of a Bohmian universe have significantly more information about $\Psi_{t_{0}}$ than what is already provided by the Past Hypothesis.
+This completes the argument.
+This conclusion covers also the possibility that observers might make observations (say, with telescopes) or experiments, as the outcomes of these observations or experiments will be recorded in the configuration of some particles (e.g., when scientists publish their findings).
+It also covers the possibility of making observations or experiments at several times $t_{1},t_{2},\ldots,t_{k}$, since the outcomes will be recorded and thus can be read off from $Q_{t}$ at $t\geq t_{k}$.
+We remark that the precise microscopic trajectories of all Bohmian particles may contain more information about $\Psi_{t_{0}}$, but it is known that observers inside a Bohmian universe cannot measure positions at several times without introducing decoherence and thereby changing the trajectory.
+Therefore, they do not have access to the full trajectories, which makes understandable why the dependence of the trajectories on $\Psi_{t_{0}}$ is compatible with OI as just derived.
+Let us summarize the second argument in a different way. A key fact in BM that leads to OI is this:
+[…]
+Corollaries 14 just provide quantitative statements (in the more general framework of POVMs) of how this near-independence makes $\Psi$ observationally indistinguishable from $\rho_{0}$ or other $\Psi$s.
+They show that that even complete information about $Q_{t}$ would not help with distinguishing between typical $\Psi_{t_{0}}$ and $\rho_{0}$ (or between two typical wave functions); a fortiori, limited information about $Q_{t}$ does not help.
+In OQM, the second argument can be applied as well (although it is a bit less clear because the ontology of OQM is less clear): after all, the Heisenberg cut (between system and apparatus) can always be moved so as to enlarge the system, and the $|\Psi_{t}|^{2}$ distribution in configuration space will also in OQM determine the distribution of macroscopic records (say, in books or journals) at time $t$.
+Thus, the macroscopic records at $t$ provide almost no information about $\Psi_{t_{0}}$.
+Concerning EQM, we have reservations about the justification of the use of probabilities (Barrett 2023, sec.
+6; Tumulka 2022b, sec.
+3.5.5); but to the extent that they can be justified, our second argument for OI applies as well.
+In GRW theory, the second argument can be applied in a slightly modified variant.
+To this end, we consider GRWf, the GRW theory with a flash ontology in spacetime (Ghirardi and Bassi 2020, sec.
+11; Tumulka 2022b, sec.
+3.3.4).
+Let $F_{t}$ denote the pattern of flashes up to $t$, with $t>t_{0}$.
+A key fact about OI in GRWf is this:
+for any subset $M\in\cup_{n=0}^{\infty}\mathbb{R}^{4n}$, the probability of the event that $F_{t}\in M$ is nearly independent of $\Psi_{t_{0}}$ for typical $\Psi_{t_{0}}\in\mathbb{S}(\mathscr{H}_{0})$.
+(18)
+This follows from Theorem 1 by taking the POVM $E$ to be the POVM governing the joint distribution of all flashes (Tumulka 2022b, sec.
+5.1.1 and 7.8).
+As for the configuration in BM, observers in a GRWf universe do not have access to the full pattern of flashes, $F$, at most to a coarse-grained version of it.
+But again, we use the fact that any record would have to be encoded in the flashes (as are the positions of pointers or the shape of printed letters).
+More precisely, we use the following
+Necessary Condition
+If an event is known to inhabitants of a GRWf universe at time $t$ then it is of the form $F_{t}\in M$ for some set $M$.
+Now Corollaries 14 show that no such event can reliably distinguish $\Psi_{A}$ from $\rho_{0}$, or can significantly favor $\Psi_{A}$ over $\rho_{0}$ or $\Psi_{B}$, or can substantially alter the Bayesian posterior distribution of $\Psi_{t_{0}}$, provided $\operatorname{tr}(\rho_{0}E(M))>\varepsilon$.
+Thus, at no time $t$ can observers in a GRWf universe have significant information about $\Psi_{t_{0}}$.
+This completes the argument.
+A difference to the Bohmian case is that the exact history of the universal configuration in BM may provide sufficient information to determine $\Psi_{t_{0}}$, whereas in GRWf even the exact history of all flashes is of no help.
+Thus, GRWf provides, in a sense, a stronger kind of limitation to knowledge than BM.
+Another difference to BM is that inhabitants of a GRWf universe living at time $t$ may very well find out a lot about the wave function $\Psi_{t}$ at their time (in contrast to the initial wave function $\Psi_{t_{0}}$).
+After all, $\Psi_{t}$ is not unitarily evolved from $\Psi_{t_{0}}$ but has collapsed; for example, if we find Schrödinger’s cat alive then its wave function will have collapsed to (approximately) the wave function of a live cat.
+(Note that this still does not give us new information about $\Psi_{t_{0}}$.)
+3.5 Remarks
+3.5.1 Experiment Framing vs $|\Psi_{t}|^{2}$ Framing
+Our first remark concerns a comparison of the two arguments given in §3.3 and §3.4.
+Although both used the Corollaries 14, they used them in different ways: the first was framed in terms of experiments, the second in terms of the $|\Psi_{t}|^{2}$ distribution in configuration space.
+The first argument pretends that the observers are somehow not part of the universe.
+Although we do not think that this approach leads to any wrong conclusions here, we note that the second argument is clear-cut and the first less so, as various questions can arise about the first: Is it not usually assumed that observers and their apparatus are not entangled with the object they are observing?
+And is that not violated in the relevant observations of the universe?
+Given that in OQM, observations can collapse wave functions, should not $E_{z}$ be applied to a collapsed wave function instead of $\Psi_{t_{0}}$?
+While none of these objections to the first argument withstands scrutiny, the second argument makes it much more transparent that the conclusion, OI, remains valid.
+The second argument also makes it clear that repeating the experiment does not allow the observer to measure the probability more precisely; after all, records of all experiments are included in the configuration at time $t$, and the whole universe cannot be repeated, of course.
+3.5.2 Other Kinds of Indistinguishability
+As already remarked, the question we ask is different from the question, given two states $\Psi_{A},\Psi_{B}$, is there a POVM that will distinguish them?
+It is also different from the question, can we empirically distinguish the possibility that the universe is in the mixed state $\rho_{0}$ from the possibility that it is in a random pure state $\Psi$ with distribution $u_{0}$?
+The answer to the latter question is trivially “no” for any value of $d_{0}$ (large or not), as any observation $z$ with POVM element $E_{z}$ has probability $\operatorname{tr}(\rho_{0}E_{z})$ in the mixed case and probability $\int_{\mathbb{S}(\mathscr{H}_{0})}u_{0}(d\psi)\langle\psi|E_{z}|\psi\rangle$ in the random pure case, and the two probabilities are always exactly equal.
+(Readers may even find it hard to see any physical difference at all between the mixed state $\rho_{0}$ and a random pure state $\Psi$ with distribution $u_{0}$; but this difference becomes visible in Bohmian mechanics, where trajectories guided by $\rho_{0}$ are different from those guided by $\Psi$.)
+Our results convey that the Born distribution $\langle\Psi|E_{z}|\Psi\rangle$ is close to $\operatorname{tr}(\rho_{0}E_{z})$ for most $\Psi$, whereas the impossibility to distinguish the mixed case from the random pure case arises from the fact that $\langle\Psi|E_{z}|\Psi\rangle$, when averaged over $\Psi$, yields $\operatorname{tr}(\rho_{0}E_{z})$.
+The difference between our results and the impossibility to distinguish the mixed case from the random pure case is perhaps most easily visible in the Bayesian statement of §3.2.3.
 4 Comparisons with Known Results
-In this section, we briefly compare our results for OI with some known results in philosophy of science and foundations of physics.
-In quantum mechanics, there are several results about limitation to knowledge about the quantum state. For a survey, see (Tumulka 2022b, ch.5). Here we highlight a few:
-  • Because of the Main Theorem about POVMs, we cannot measure the wave function of a given system.
+We compare observation typicality with some known results in foundations of physics.
+There are several results about limitation to knowledge about the quantum state.
+For a recent survey, see (Tumulka 2022b, ch.5).
+Here we highlight a few: (1) Because of the Main Theorem about POVMs, we cannot measure the wave function of a given system with useful precision.
 Relatedly, because of the no-cloning theorem, we cannot reliably copy the wave function of a given system.
-  • It can be shown (Tumulka 2022a) that in any ontological theory of quantum mechanics, it is impossible to measure the ontic state.
+(2) It can be shown (Tumulka 2022a) that in any ontological theory of quantum mechanics, it is impossible to measure the ontic state.
 Thus, whatever the true theory of quantum mechanics may be, there must be facts that cannot be empirically determined.
-  • Because of the Pusey-Barrett-Rudolph (PBR) theorem (Pusey et al.
+(3) Because of the Pusey-Barrett-Rudolph (PBR) theorem (Pusey et al.
 2012), two different ensembles of wave functions with the same density matrix are physically distinct.
@@
 Moreover, all three items are compatible with substantive learning about the actual quantum state.
-One may still hope to rule out a significant fraction of quantum states based on some observation.
-Moreover, a Bayesian can, in general, update their probability distribution based on observations after which their posterior may be highly peaked.
-In contrast, observation typicality tells us that observations in a typical quantum universe do not rule out more than a tiny fraction of universal quantum states, and (if the observation is not too improbable) do not have a substantial influence on a uniform prior probability distribution about initial quantum states.
-It stays nearly uniform.
-We now compare our results to a much discussed underdetermination result in the philosophy of general relativity (GR).
-Manchak (2009) showed that for almost every spacetime model of GR, there exists another spacetime that is physically distinct (in the sense of having a different global structure) but observationally indistinguishable from it.
+One may still have hoped to reliably eliminate a significant fraction of quantum states based on some observation.
+One might also have hoped that, a Bayesian could, in general, update their probability distribution based on observations, leading to a highly peaked posterior.
+However, observation typicality conveys that these are impossible for the universal quantum state.
+Observations in a typical quantum universe do not reliably eliminate more than a tiny fraction of universal quantum states, and—if the observation is not too improbable—do not significantly reshape a uniform prior probability distribution over initial quantum states.
+The distribution remains nearly uniform.
+In the philosophy of general relativity (GR), Manchak (2009, 2011) showed that for every spacetime model of GR of a certain class, there exists another spacetime that is physically distinct (in the sense of having a different global structure) but observationally indistinguishable from it.
 [^17] For some philosophical discussions about this result, see Beisbart (2009), Norton (2011), Butterfield (2012), and Cinti and Fano (2021).
-Both Manchak’s result and ours are compatible with a number of inductive hypotheses.
-In Manchak’s case, one can assume that the same law of nature (Einstein’s equation of GR) applies to all spacetime points.
-Moreover, Manchak’s result is compatible with any set of local conditions (conserved under local isometries).
-In our case, one can also assume that the same laws of nature (within quantum theory) apply to the entire model, so that we are considering the same Hilbert space $\mathscr{H}$, the same Hamiltonian $H$, and the same subspace $\mathscr{H}_{0}$.
-There are important differences too.
-First, Manchak’s argument shows the existence of an OI counterpart for almost every model of GR, while our argument shows that almost every model is an OI counterpart to almost every other model of QM.
+Remarkably, both GR and QM, the two pillars of modern physics, imply OI.
+We highlight two differences between the arguments.
+First, Manchak’s argument shows the existence of an OI counterpart for every model of GR in a special class.
 Hence, his result is compatible with the fact that we can observationally rule out a significant fraction of possible universes while ours is not.
 Second, Manchak’s proof uses a “cut-and-paste” method to construct special models of GR that are OI from the original model, while ours uses the probabilistic method to show that OI must hold among typical models of QM with generic features.
-Related to the second point, Manchak’s argument has been challenged on philosophical grounds that the constructed spacetimes are not “physically reasonable,” because some people regard the “cut-and-paste” models as implausible candidates for the actual universe.
-(For example, see Cinti and Fano (2021).) One may respond, as Manchak does, that “the theorem can have physical relevance even if the particular model constructed in the proof does not” (Manchak 2011, p.7).
-We do not take a stance on that issue, but merely point out that the objection of “physical unreasonableness” does not even work as a prima facie objection against our argument for OI in QM.
+Related to this point, Manchak’s argument has been challenged on philosophical grounds that the constructed spacetimes are not “physically reasonable,” because some people regard the “cut-and-paste” models as implausible candidates for the actual universe.
+(For example, see Cinti and Fano (2021).) We do not take a stance on that issue, but merely point out that the objection of “physical unreasonableness” does not even work as a prima facie objection against our argument for OI in QM.
 We do not show that almost any model has at least one specially constructed counterpart.
 Rather, we show that typical (generic) models of QM in a high-dimensional Hilbert space are OI from each other.
-It is interesting to note that GR and QM, two pillars of modern physics, both lead to OI.
 5 Philosophical Implications
-We highlight some potential implications of observation typicality for philosophy of science.
-First, observation typicality places severe limitations to knowledge in a quantum universe.
-In a large universe, no experiment can distinguish the overwhelming majority of universal quantum states.
-This is an in-principle limitation that cannot be overcome by new technologies.
-Second, the flip side is that, concerning the observable properties of the universal quantum state, we may know a lot about them because of the typicality results.
-For any observation, almost all universal quantum states look the same as the normalized projection $\rho_{0}$.
-Assuming our universe is typical among those that satisfy the PH, even though we do not know which it is exactly, we know there is uniformity in their predictions—they are very nearly the same as the probabilistic prediction of the Wentaculus density matrix.
-Fixing the PH nearly fixes the probabilistic predictions of typical quantum states.
-It provides a new perspective on the predictive power of initial condition laws such as the PH.
-When the initial condition law delivers a high-dimensional subspace in the full Hilbert space, the precise choice of a typical quantum state makes almost no difference to the probabilistic prediction.
-For any observation, typical individual quantum states predict with nearly the same probability distribution.
-This raises an interesting question, which we leave to future work, whether typical choices of the initial condition, in some sense to be defined (e.g.
-typical subspace), lead to nearly the same probability distributions.
+We sketch four potential philosophical implications of observation typicality. Our aim here is exploratory, not prescriptive.
+First, observation typicality imposes severe limitations on knowledge in a quantum universe.
+They are in-principle limitations that cannot be overcome, regardless of technological advancements, except in cases where the relevant provisos do not apply.
+Second, the flip side is that observation typicality allows us to know a great deal about the observable properties of the universal quantum state.
+For any given observation, nearly all universal quantum states are indistinguishable from the normalized projection $\rho_{0}$.
+Hence, fixing the Past Hypothesis almost completely determines the probabilistic predictions of typical quantum states in $\mathscr{H}_{0}$.
 Third, observation typicality may conflict with positivism, the idea that a statement is unscientific or meaningless if it cannot be tested experimentally, and a variable is not well-defined if it cannot be measured.
-Given the PBR theorem, and the various solutions to the quantum measurement problem, we have good reasons to accept that the universal quantum state is an objective feature of the universe.
-Due to observation typicality, however, it is undetertermined by observational evidence.
-This is ironic, to echo a point made by Cowan and Tumulka (2016): positivism is sometimes supported by examples from quantum mechanics, and yet quantum mechanics tells us that the central object in a quantum universe is well-defined, objective, but non-measurable, which undermines positivism.
-Finally, our result has potential implications for realism about the quantum state.
-On realism, it is an open question what the universal quantum state represents and how best to understand its objectivity and reality.
-(For a survey, see Chen (2019).) Within realism, there are two main approaches.
-The first is to regard it as a physical thing, like a physical field, that exists as part of the basic building blocks of the universe (Albert 1996, Ney 2021).
-The second is to regard it as a physical law, like the Hamiltonian function in classical mechanics, that tells physical things (such as particles and other fields) how to move (Goldstein and Zanghì 2013).
-If the quantum state is a physical thing, OI suggests an in-principle limit to how precisely it can be measured: we cannot gain substantive information about which quantum state is the actual one.
-Physical objects in classical mechanics do not have this character.
-They can be measured to arbitrary precision.
-Almost all quantum states of the universe will appear the same to us.
-This should be challenging to the empiricists who hold out hope that the basic building blocks of the universe should be in-principle measurable.
-Perhaps the result is less surprising if the universal quantum state has the status of a physical law.
-Since it is underdetermined by observational evidence, we have to use theoretical virtues, such as simplicity and elegance, to pin it down.
-The theoretical flavor of how we choose the universal quantum state is similar to how we choose other physical laws, whose precise forms are constrained but still vastly underdetermined by observational evidence.
-Realism about the universal quantum state also leaves open whether it is necessarily a pure state represented by a wave function, or possibly an impure (“mixed”) one represented by a density matrix.
-Following Chen (2021, 2024b), we call the first option wave function realism (WFR) and the second density matrix realism (DMR).
-It is already known that WFR and DMR are empirically equivalent, given appropriate choices of the universal quantum states.
-Observation typicality suggests that it is even harder to distinguish the two theories.
-Not only are typical individual wave functions OI from each other, but they are also OI from the normalized projection, an impure density matrix.
-One may regard it as a potential reason to prefer DMR over WFR.
-If typical pure states $\Psi_{t_{0}}$ compatible with the Past Hypothesis do not lead to different predictions, why not just use the Wentaculus density matrix $\rho_{0}$ instead?
-After all, $\rho_{0}$ may be regarded as much simpler than any typical $\Psi_{t_{0}}$ even though they are nearly predictively equivalent.
-Here we do not take a firm stance on these implications, merely exploring how they may be suggested by our results.
+Given the PBR theorem (Pusey et al.
+2012), and the various solutions to the quantum measurement problem, we have good reasons to accept that the universal quantum state $\Psi$ is an objective feature of a quantum universe, yet it is observationally underdetermined because of observation typicality.
+This is ironic, to echo a point made by Cowan and Tumulka (2016): positivism is sometimes defended using examples from quantum mechanics, yet quantum mechanics itself reveals that a central object in a quantum universe—the universal quantum state—is well-defined, objective, but fundamentally unmeasurable.
+Finally, our results impact quantum state realism (see Chen (2019) for a survey).
+One view treats the universal quantum state as a physical entity—a field constituting one of the universe’s basic building blocks (Albert 1996, Ney 2021).
+Yet observation typicality implies limits to even imprecisely measuring it, challenging the empiricist expectation that fundamental constituents of the universe are in principle measurable (at least imprecisely).
+Alternatively, if we interpret the universal quantum state as a physical law (akin to the classical Hamiltonian; see Goldstein and Zanghì (2013)), we may naturally expect it to be constrained by observationaly evidence but remain vastly underdetermined, much like other physical laws.
+Quantum state realism leaves open the question of whether it must necessarily be a pure state, represented by a wave function, or if it could be a mixed state, represented by a density matrix.
+Following Chen (2021, 2024b), we refer to the first option as wave function realism (WFR) and the second as density matrix realism (DMR).
+Not only are typical individual wave functions OI from each other, but they are also OI from $\rho_{0}$, which is mixed.
+This may offer a defeasible reason to prefer DMR over WFR.
+If typical pure states $\Psi_{t_{0}}$ compatible with the Past Hypothesis yield nearly identical predictions as $\rho_{0}$, why not adopt the Wentaculus version of DMR and postulate the Wentaculus density matrix $\rho_{0}$ instead?
+After all, $\rho_{0}$ can be regarded as much simpler than any typical $\Psi_{t_{0}}$ while maintaining near predictive equivalence.
 6 Conclusion
-We have shown that the central object in quantum theory—the quantum state of the universe—is most likely hidden from observation, because typical quantum states in a high-dimensional Hilbert subspace $\mathscr{H}_{0}$ are observationally indistinguishable from the density matrix $\rho_{0}=P_{0}/d_{0}$ and thus from each other.
-In fact, no observation will yield any substantial information at all, even partial or probabilistic, about $\Psi_{t_{0}}$.
-That is in a sense the strongest known result about a limitation to knowledge in a quantum universe.
-Assuming the universe we inhabit is typical, our observational data alone will tell us very little about exactly which one it is that we inhabit.
-Nature is more secretive than we have realized.
+We have shown that a central object in quantum theory—the quantum state of the universe—is effectively hidden from observation if it is typical.
+Typical quantum states in a high-dimensional Hilbert subspace $\mathscr{H}_{0}$ are observationally indistinguishable from the density matrix $\rho_{0}=P_{0}/d_{0}$ and, consequently, from each other.
+No observation will yield substantial information about $\Psi_{t_{0}}$—not reliably and (unless the observation is extremely unlikely) not even unreliably.
+These results amount to the strongest epistemic limits we know for a quantum universe.
+If the wave function of our own universe is typical in $\mathscr{H}_{0}$, our observational data alone will reveal very little about which quantum state it is in.
+Nature, it seems, is far more secretive than we have realized.
 Acknowledgment.
-We thank Sheldon Goldstein for extensive feedback and for suggesting the term “observation typicality.” For helpful discussions, we also thank Emily Adlam, David Albert, Fabio Anza, Patricio Avila Cardenas, Jeff Barrett, Jeremy Butterfield, Craig Callender, Branden Fitelson, Hans Halvorson, Alan Hájek, Tyler Hildebrand, Christopher Jarzynski, Joel Lebowitz, JB Manchak, Kerry McKenzie, Marcos Rigol, Daniel Rubio, Tony Short, Stefan Teufel, Isaac Wilhelm, Nicole Yunger Halpern, the UC San Diego Philosophy of Physics Group, and the participants at the 2024 Quantum Thermodynamics Workshop at the University of Maryland, College Park.
+We thank Sheldon Goldstein for his extensive feedback and for suggesting the term “observation typicality.” For helpful discussions, we are also grateful to Emily Adlam, David Albert, Fabio Anza, Patricio Avila Cardenas, Jeff Barrett, Jeremy Butterfield, Craig Callender, Jennifer Carr, Silvia De Bianchi, Branden Fitelson, Hans Halvorson, Alan Hájek, Tyler Hildebrand, Nick Huggett, Christopher Jarzynski, Andrew Jordan, Joel Lebowitz, Matthew Leifer, Barry Loewer, JB Manchak, Tim Maudlin, Travis McKenna, Kerry McKenzie, Marcos Rigol, Daniel Rubio, Charles Sebens, Shelly Yiran Shi, Tony Short, Stefan Teufel, Mordecai Waegell, Isaac Wilhelm, Nicole Yunger Halpern, the CalTech Philosophy of Science Group, the SoCal Quantum Foundations Hub, the Metro Area Philosophy of Science Group, the COSMOS research team, the UC San Diego Philosophy of Physics Group, the participants at the 2024 Quantum Thermodynamics Workshop at the University of Maryland, College Park, and the 2025 APS March Meeting.
 Funding.